### v1

# Certified spectral approximation of transfer operators and the Gauss map

Abstract
We prove that the full spectral picture of a transfer operator (every isolated eigenvalue, eigenvector, and Riesz projector outside the essential spectral radius) can be approximated to arbitrary precision by finite-rank discretizations, with no spectral pollution. The method is a posteriori: once a computable approximation bound is available (from compactness or a Doeblin–Fortet–Lasota–Yorke inequality, which may require hyperbolicity constants and adapted Banach spaces), the spectral gap and multiplicity are certified from computed data via a resolvent perturbation bound, with no further dynamical input. This applies both to compact operators on a single Banach space and to quasi-compact operators satisfying a Doeblin–Fortet–Lasota–Yorke inequality, extending Li’s resolution of the Ulam conjecture from the invariant density to the entire discrete spectrum with certified error bounds at every finite truncation.
As a benchmark, we certify the first $50$ nonzero eigenvalues of the Gauss–Kuzmin–Wirsing operator to at least $90$ rigorous decimal digits, together with their eigenvectors, Riesz projectors, and spectral gap, yielding a certified spectral expansion for Gauss–Kuzmin distributions with explicit error bounds and providing a rigorous answer to the Gauss–Babenko–Knuth problem on the spectral data of the Gauss map.
MSC 2020: 37C30, 47A75 (primary); 37M25, 65G20, 11K50, 47A58 (secondary).
1 Introduction and overview
Spectral questions as certified linear algebra.
Transfer operators encode invariant measures, rates of mixing, limit theorems, and Ruelle–Pollicott resonances, yet extracting quantitative spectral information from an infinite-dimensional operator remains delicate: one must control discretization/truncation effects, resolvent growth, and spectral-projector stability with fully explicit constants.
The central message of this paper is that, under broadly available regularization hypotheses (compactness/nuclearity on a suitable space, or a Lasota–Yorke inequality), the spectral analysis of an infinite-dimensional operator can be reduced to a finite-dimensional, computationally certifiable task. Starting from a finite-rank discretization $L_{N}$ of an operator $L$ with an explicit error bound, we compute spectral objects for $L_{N}$ (eigenvalues, resolvent bounds, Riesz projectors) and transfer them to $L$ via a resolvent perturbation bound with fully explicit constants. The analytic work is confined to producing a few quantitative constants, after which the spectral certification becomes validated numerical linear algebra. The a posteriori approach to certifying spectral data of transfer operators from computed information originates in [28, 27, 29, 26, 25], which developed rigorous certification of invariant measures, mixing rates, and escape rates via Lasota–Yorke inequalities and the coarse-fine approximation strategy. [12] extended the approach to eigenvalue enclosures and Riesz projector certification via validated linear algebra. Here we isolate a set of hypotheses applicable in a range of functional-analytic settings, extend the framework to the Doeblin–Fortet–Lasota–Yorke (DFLY) setting, and develop high-precision implementations that certify many eigenvalues, eigenvectors, and Riesz projectors simultaneously.
Over the past two decades, anisotropic Banach spaces and Lasota–Yorke inequalities have been constructed for broad classes of hyperbolic and piecewise hyperbolic systems [11, 32, 24, 5], yet for many concrete models (higher-dimensional piecewise hyperbolic maps, billiards with complex geometry, perturbations of known systems [23, 15]), proving a spectral gap on these spaces remains a difficult open problem. The present framework reduces the spectral gap question for any such system to a finite computation, producing certified eigenvalue enclosures and mixing rates. Even where a spectral gap has been established analytically, the method provides quantitative data (the actual gap size, subdominant eigenvalues, and projector norms) that purely qualitative arguments do not. Crimmins and Froyland [21] developed Fourier-analytic discretizations on the Gouëzel–Liverani spaces and proved spectral convergence for Anosov maps on tori, but their framework assumes the spectral gap a priori (the operator is “simple quasi-compact” in the Keller–Liverani sense) and establishes convergence of the approximations without explicit error bounds at finite truncation. The present work removes both limitations: the spectral gap itself is certified a posteriori from computed data, with explicit error bounds at any given matrix size.
The Gauss–Kuzmin–Wirsing problem.
As a benchmark, we consider one of the oldest problems in ergodic theory and thermodynamic formalism: the Gauss map $T(x)=\{1/x\}$ on $(0,1]$, which generates continued fractions. Let $x_{0}$ be uniformly distributed on $[0,1]$, and set $x_{n}:=T^{n}(x_{0})$. Gauss [30] (English translation in [53, Appendix III]) asked for the distribution function
$F_{n}(x)=\mathbb{P}(x_{n}\leq x),$
whose limit $F_{\infty}(x)=\log_{2}(1+x)$ corresponds to the invariant probability measure of $T$. This distribution is governed by iterates of the transfer operator
$(\mathcal{L}f)(x)=\sum_{k\geq 1}\frac{1}{(x+k)^{2}}\,f\!\left(\frac{1}{x+k}\right),\qquad F_{n}(x)=\int_{0}^{x}(\mathcal{L}^{n}\mathbf{1})(t)\,dt.$
The Gauss–Kuzmin–Lévy problem concerns the rate at which $F_{n}$ converges to $F_{\infty}$. Early quantitative bounds go back to Kuz’min [39] and Lévy [41], while Wirsing [56] identified the leading exponential rate via the subdominant spectrum of the Gauss–Kuzmin–Wirsing (GKW) operator. See also Mayer [44, 45] and the eigenvalue asymptotics studied by Alkauskas [1].
The problem of quantifying this rate already appears as Exercise 22 [HM50] in the first edition of Knuth [37, p. 337]:
(K. F. Gauss.) What is the asymptotic behavior of $F_{n}(x)-\log_{2}(1+x)$ as $n\to\infty$?
This predates both the operator-theoretic resolution of Wirsing [56] and the discretization-based approaches of Babenko [4, 3]. In the third edition [38, p. 376], the exercise is reformulated in spectral terms, reclassified to [HM46], and attributed to Babenko:
(K. I. Babenko.) Develop efficient means to calculate accurate approximations to the quantities $\lambda_{j}$ and $\Psi_{j}(x)$, for small $j\geq 3$ and for $0\leq x\leq 1$.
The shift from Gauss’s distributional question to Babenko’s spectral formulation reflects the mathematical progress between editions.
The GKW operator has been the subject of extensive high-precision computations; however, as emphasized in [1], rigorous certification has historically been available only for the first few digits of the leading subdominant eigenvalues.
Certified spectral data for the Gauss map.
Theorem 1.1.
The first $50$ nonzero eigenvalues of the Gauss–Kuzmin–Wirsing operator, realized as $L:=S\mathcal{L}:H^{2}(D_{1})\to H^{2}(D_{1})$ (Section 4), are real and simple. Certified enclosures are given in the supplementary material, each with at least $90$ certified decimal digits. In particular, the subdominant eigenvalue (the Gauss–Kuzmin–Wirsing constant) satisfies $|\lambda_{2}-\tilde{\lambda}_{2}|<10^{-175}$, where
[…]
$23787\,91324\,46139\,89647\,64297.$
Certified enclosures for the corresponding eigenvectors and Riesz spectral projectors are provided, yielding a validated spectral expansion
$L^{n}\mathbf{1}=\sum_{j=1}^{50}\lambda_{j}^{n}\,\ell_{j}(\mathbf{1})\,v_{j}+R_{50}(n),\qquad\|R_{50}(n)\|_{H^{2}(D_{1})}\leq 3.72\times 10^{21}\cdot(1.01\times 10^{-21})^{n+1},$
for the Gauss–Kuzmin distributions, with rigorous error below $10^{-20}$ already at $n=1$.
Theorem 1.1 provides rigorous certified solutions to the Gauss–Babenko–Knuth problem: we compute $\lambda_{j}$ and $\Psi_{j}$ for $j=1,\ldots,50$ with at least $90$ validated digits each, together with the spectral projectors needed for the expansion of $F_{n}$. Both the number of eigenvalues and the precision can be pushed further: the truncation error decays geometrically in the matrix size, so additional digits or eigenvalues require only larger matrices and higher-precision arithmetic, with no new analytic input.
Figure 1: The first six certified eigenfunctions $v_{1},\ldots,v_{6}$ of the GKW transfer operator on $[0,1]$, computed at $K=512$ (1024-bit precision). Each panel shows the eigenvalue $\lambda_{j}$ and spectral coefficient $\ell_{j}(\mathbf{1})$. The invariant density $v_{1}=\frac{1}{\ln 2}\cdot\frac{1}{1+x}$ (top left) is positive; higher eigenfunctions develop a singularity at $x=0$ and oscillations that reflect the Markov partition of the Gauss map.
Spectral gap as an output, not an input.
Spectral separation is certified a posteriori from computed data, rather than assumed a priori. Many rigorous computational frameworks in thermodynamic formalism require a qualitative spectral gap as a starting hypothesis, from which quantitative consequences are then derived [48, 49, 55]. The present method certifies spectral gaps directly: the resolvent bounds along contours that enclose individual eigenvalues simultaneously establish spectral separation, eigenvalue multiplicity, and projector stability. This provides the certified spectral data that such approaches take as input. For Markov operators, this has a direct probabilistic interpretation: the spectral gap governs the mixing rate of the associated chain, so a posteriori spectral certification turns mixing rates into certified outputs rather than assumed inputs.
The general framework.
Although the GKW operator is the main benchmark, the method applies whenever the following inputs are available for a bounded operator $L:\mathcal{B}\to\mathcal{B}$ on a Banach space: (A) a finite-rank approximation $L_{N}$ with an explicit error bound $\|L-L_{N}\|\leq\varepsilon_{N}$; (B) certified resolvent bounds $\|(zI-L_{N})^{-1}\|$ along contours $\Gamma\subset\mathbb{C}$, obtained by validated linear algebra (Schur/SVD, interval/ball arithmetic); and (C) a resolvent perturbation bound transferring certified bounds from $L_{N}$ to $L$. The key condition is $\varepsilon_{N}\sup_{z\in\Gamma}\|(zI-L_{N})^{-1}\|<1$; when it holds, spectral enclosures, Riesz projectors, and invariant subspaces of $L_{N}$ transfer to $L$ with explicit error control. The full framework is developed in Section 2; holomorphic transfer operators provide a natural setting where (A) arises from compactness/nuclearity induced by holomorphic extension, and related analytic frameworks appear in, e.g., [7, 48]. For the Gauss map, the holomorphic contraction of the branches induces compactness on Hardy spaces with explicit truncation bounds (input (A) with exponential decay $(2/3)^{K}$), and certified Schur decompositions provide the resolvent control (input (B)). In an appendix we show that the same philosophy extends to the strong-weak setting of Doeblin–Fortet–Lasota–Yorke (DFLY) inequalities, where the operator $L$ acts on a Banach space $\mathcal{B}_{s}$ equipped with a weaker norm $\|\cdot\|_{w}$ and satisfies a Lasota–Yorke inequality $\|Lf\|_{s}\leq a\|f\|_{s}+b\|f\|_{w}$. Input (A) becomes the strong-to-weak approximation bound $\|L-L_{k}\|_{s\to w}\leq\delta_{k}$, and the resolvent of $L_{k}$ is first controlled in the weak norm, which is computable from a finite matrix, and then lifted to the strong norm via the Lasota–Yorke inequality.
Guaranteed convergence of the full spectral picture.
A second main contribution, beyond the GKW application, is the following convergence guarantee.
Theorem 1.2.
Let $L$ be a bounded operator on a Banach space, and let $(L_{K})$ be a sequence of finite-rank discretizations with $\|L-L_{K}\|\leq\varepsilon_{K}\to 0$ (compact case, Theorem 2.15) or $\|L-L_{k}\|_{s\to w}\leq\delta_{k}\to 0$ under a DFLY inequality, subject to a rate condition linking $\delta_{k}$ to the norm-equivalence constants of the discretization spaces (strong-weak case, Theorem A.18). Let $U\subset\mathbb{C}$ be a spectral window isolating finitely many eigenvalues of $L$ outside the essential spectral radius. Then for $K$ large enough:
  1. 1.
    (No spectral pollution.) $L_{K}$ and $L$ have the same number of eigenvalues (counted with multiplicity) in $U$.
  2. 2.
    (Projector convergence.) $\|P_{L_{K}}(U)-P_{L}(U)\|\to 0$ as $K\to\infty$.
  3. 3.
    (A posteriori certification.) The perturbation condition that triggers (1)–(2) is checkable from the computed data of $L_{K}$ alone, with no spectral gap assumption.
In the compact case, every isolated nonzero eigenvalue, eigenvector, and Riesz projector of $L$ can be approximated to arbitrary precision by the finite-rank data of $L_{K}$. In the DFLY case, the same holds for eigenvalues whose modulus exceeds any fixed $\mu>a$, provided the rate condition is satisfied at $\mu$.
These results are stated in abstract operator-theoretic terms. For quasi-compact operators satisfying a DFLY inequality, the spectrum outside the essential spectral radius consists of isolated eigenvalues of finite multiplicity, and the discrete spectral data there can be recovered from finite-rank discretizations, subject to a rate condition on the discretization error. Theorem A.18 makes this precise and quantitative. Li [42] proved convergence of Ulam approximations to the invariant density for piecewise expanding maps; the Keller–Liverani perturbation theorem [36] extends this to the general DFLY framework (see, e.g., [43, 25]) and implies that isolated eigenvalues outside the essential spectral radius are stable under discretization, qualitatively extending convergence from the invariant density to the full discrete spectrum. The present work makes this extension certified and quantitative: the Galerkin hierarchy $L_{k}$ approximates all eigenvalues, eigenvectors, and spectral projectors outside the essential spectral radius, with explicit error bounds at every finite truncation.
The Keller–Liverani spectral stability theorem [36] and related frameworks [23, 15] assume a priori that certain eigenvalues are isolated and deduce their stability under perturbation. Liverani [43] reversed the logical direction by proposing to use the finite-rank (Ulam) approximation as the “known” operator and certify spectral data of the original operator a posteriori, via the Keller–Liverani perturbation theory; this was presented as a feasibility study without rigorous computations, and the required strong resolvent norms of the discretization are nontrivial to compute in practice. The present work addresses these difficulties with the a posteriori certification tools described above. In both settings the computable truncation bound is of dynamical origin: in the compact case it arises from the regularity of the underlying map (nuclearity, domain gain on Hardy spaces), while in the DFLY case it requires hyperbolicity constants, adapted Banach spaces, and a Lasota–Yorke inequality. Once these bounds are in hand, the entire spectral certification is reduced to a computer program whose output is a rigorous proof.
Computability of isolated spectral data.
A longstanding open problem, raised by Arveson [2] and Böttcher [13], asks whether the spectrum of a bounded operator can be computed from its finite-dimensional sections. The Solvability Complexity Index (SCI) hierarchy developed by Colbrook, Hansen, and collaborators [10, 17, 18] resolves this question by showing that, for general bounded operators on a separable Hilbert space, computing the spectrum requires several successive limits and no single algorithm can provide error control at any finite step [10, Theorem 1.1]. Even for compact operators the obstruction persists: convergent algorithms exist, but no algorithm that takes only matrix entries as input can certify that its output is within a prescribed distance of the true spectrum.
The a posteriori framework of Theorems 2.15 and A.18 resolves the problem in the affirmative for the discrete spectrum of transfer operators with explicit truncation bounds. For each isolated eigenvalue, the certification algorithm takes a single limit ($K\to\infty$) and provides verified error bounds at every truncation where the $z$-dependent perturbation condition $\varepsilon_{K}\,\mathcal{R}(z,L_{K})<1$ is satisfied along the enclosing contour. The error bound at each eigenvalue depends on $z$ through the resolvent of the finite-rank model $L_{K}$, so certification targets individual eigenvalues via contour-specific resolvent control rather than uniform approximation of the full spectrum. The input that circumvents the general impossibility is the computable truncation bound $\varepsilon_{K}$, which is not available for abstract compact operators given only by matrix entries but is furnished by the dynamical structure of the transfer-operator setting. Since every compact operator with positive spectral radius is quasi-compact (its essential spectral radius is zero), the impossibility results for general compact operators suggest that the Lasota–Yorke inequality is a strictly stronger condition than quasi-compactness in a computational sense: it provides precisely the oracle enrichment needed to break the SCI barrier for spectral certification, and the impossibility results indicate that this structure cannot be universally available or uniformly recoverable from finite-dimensional data.
Related approaches.
A substantial literature develops rigorous computer-assisted estimates for transfer operators. Representative examples include explicit eigenvalue bounds on holomorphic function spaces [6, 7], rigorous schemes based on Lagrange–Chebyshev approximation and min-max techniques [52, 48], and thermodynamic computations for Hausdorff dimension problems [49, 55]. As noted above, such methods typically assume spectral separation as input; the a posteriori approach developed here and in [28, 27, 29, 26, 25, 12] provides precisely that input, suggesting a hybrid strategy: a posteriori certification via coarse (e.g. Ulam) discretizations can establish spectral gaps and isolation, after which exponentially convergent methods (e.g. on holomorphic function spaces) can be deployed for high-precision computation. In the single-space (compact) setting, this can even bootstrap: a crude initial bound on the spectral gap can be sharpened a posteriori to arbitrary precision, as we demonstrate for the Gauss map. In the two-space (DFLY) setting, however, a priori methods that assume a spectral gap often cannot certify it from their own output, making the a posteriori approach essential in many cases.
A parallel program addresses the rigorous computation of spectra and pseudospectra for general classes of operators, including verified algorithms for spectral inclusion sets [17, 16] and data-driven spectral methods with convergence guarantees for Koopman and transfer operators [19, 14]. These approaches typically work in Hilbert-space ($L^{2}$ or $\ell^{2}$) settings with operators specified by matrix entries or trajectory data, whereas the present method exploits explicit dynamical structure (regularity and domain gain, DFLY inequalities) for operator-norm certification of isolated spectral data with arbitrary precision.
Organization.
Section 2: abstract certified-spectral template. Section 3: certified resolvent bounds via validated Schur decompositions. Section 4: Hardy-space framework and truncation bounds for the Gauss operator. Section 5: certified spectral enclosures and eigendata. Appendix: DFLY resolvent lifting.
2 A posteriori validated spectral certification: an abstract template
The resolvent perturbation bound at the core of this section can be viewed as a quantitative, a posteriori counterpart of the Keller–Liverani spectral stability theorem [36]: rather than deducing spectral convergence from an abstract approximation scheme, we extract explicit enclosures from validated resolvent data. We work in the simplest setting: a compact operator on a single Banach space $\mathcal{B}$. The strong-weak (DFLY) framework is treated in Appendix A.
Assumption 2.1.
Let $\mathcal{B}$ be a complex Banach space and let $L\in\mathcal{L}(\mathcal{B})$ be compact. Assume we are given a sequence of finite-rank operators $\pi_{K}\in\mathcal{L}(\mathcal{B})$ and define
$L_{K}:=\pi_{K}\,L\,\pi_{K}\qquad(K\geq 1).$
We assume:
  1. 1.
    Each $\pi_{K}$ is a projection with $\pi_{K}^{2}=\pi_{K}$, $\|\pi_{K}\|\leq 1$, and the family is nested: for all $K^{\prime}\geq K$,
    $\pi_{K}\pi_{K^{\prime}}=\pi_{K}.$
    or equivalently, $\operatorname{Ran}\pi_{K}\subset\operatorname{Ran}\pi_{K^{\prime}}$.
  2. 2.
    There exists $C\geq 0$ such that
    $\|L\|_{\mathcal{B}\to\mathcal{B}}\leq C,\qquad\|L_{K}\|_{\mathcal{B}\to\mathcal{B}}\leq C\quad\text{for all }K\geq 1.$
  3. 3.
    There exists a computable sequence $\varepsilon_{K}\downarrow 0$ such that
    $\|(\mathrm{Id}-\pi_{K})L\|_{\mathcal{B}\to\mathcal{B}}\;+\;\|L(\mathrm{Id}-\pi_{K})\|_{\mathcal{B}\to\mathcal{B}}\ \leq\ \varepsilon_{K}\quad\text{for all }K\geq 1.$
    In particular,
    $\|L-L_{K}\|_{\mathcal{B}\to\mathcal{B}}\leq\varepsilon_{K},\qquad\|L_{K^{\prime}}-L_{K}\|_{\mathcal{B}\to\mathcal{B}}\leq\varepsilon_{K}\quad\text{for all }K^{\prime}\geq K.$
2.1 Resolvent inequalities and the Householder argument
Throughout, $\|\cdot\|$ denotes the operator norm on $\mathcal{B}$. For $A\in\mathcal{L}(\mathcal{B})$ we write $R_{A}(z):=(z\,\mathrm{Id}-A)^{-1}$ for the resolvent ($z\in\rho(A)$), $\mathcal{R}(z,A):=\|R_{A}(z)\|$ for the resolvent norm (with $\mathcal{R}(z,A):=+\infty$ for $z\in\sigma(A)$), and $\Lambda_{\varepsilon}(A):=\{z\in\mathbb{C}:\mathcal{R}(z,A)\geq 1/\varepsilon\}$ for the $\varepsilon$-pseudospectrum.
Lemma 2.2 ([35, Ch. IV, §1.1]).
Let $A,B\in\mathcal{L}(\mathcal{B})$ and $z\in\rho(A)\cap\rho(B)$. Then
$R_{A}(z)-R_{B}(z)=R_{A}(z)\,(A-B)\,R_{B}(z),$
and consequently
$\|R_{A}(z)-R_{B}(z)\|\ \leq\ \mathcal{R}(z,A)\,\|A-B\|\,\mathcal{R}(z,B).$
Corollary 2.3.
Let $A,B\in\mathcal{L}(\mathcal{B})$ and $z\in\rho(B)$. If
$\alpha(z):=\|(A-B)R_{B}(z)\|<1,$
then $z\in\rho(A)$ and
$\mathcal{R}(z,A)\leq\frac{\mathcal{R}(z,B)}{1-\alpha(z)},\qquad\|R_{A}(z)-R_{B}(z)\|\leq\frac{\mathcal{R}(z,B)^{2}\,\|A-B\|}{1-\alpha(z)}.$
Proof.
Factor
$z\,\mathrm{Id}-A=\bigl(\mathrm{Id}-(A-B)R_{B}(z)\bigr)\,(z\,\mathrm{Id}-B),$
invert the first term by Neumann series, and use Lemma 2.2. ∎
Proposition 2.4.
Let $L\in\mathcal{L}(\mathcal{B})$ be compact and let $L_{K}$ be finite rank with
$\|L-L_{K}\|\leq\varepsilon_{K}.$
Then
$\sigma(L)\ \subset\ \Lambda_{\varepsilon_{K}}(L_{K}).$
Moreover, let $U$ be a bounded connected component of $\Lambda_{\varepsilon_{K}}(L_{K})$ with $0\notin\overline{U}$, and assume $\partial U$ is a piecewise $C^{1}$ Jordan curve. Then $\partial U\subset\rho(L_{K})$ and
$\sup_{z\in\partial U}\ \varepsilon_{K}\,\mathcal{R}(z,L_{K})\ <\ 1,$
so $\partial U\subset\rho(L)$, and $L$ and $L_{K}$ have the same total algebraic multiplicity of eigenvalues in $U$:
$\dim\operatorname{Ran}P_{L}(U)\ =\ \dim\operatorname{Ran}P_{L_{K}}(U),\qquad P_{A}(U):=\frac{1}{2\pi i}\int_{\partial U}(zI-A)^{-1}\,dz.$
Proof.
Let $\lambda\in\sigma(L)$. If $\lambda\notin\Lambda_{\varepsilon_{K}}(L_{K})$, then $\lambda\in\rho(L_{K})$ and $\mathcal{R}(\lambda,L_{K})<1/\varepsilon_{K}$, hence
$\|(L-L_{K})(\lambda I-L_{K})^{-1}\|\leq\varepsilon_{K}\,\mathcal{R}(\lambda,L_{K})<1.$
Thus $\lambda I-L=\bigl(I-(L-L_{K})(\lambda I-L_{K})^{-1}\bigr)(\lambda I-L_{K})$ is invertible, contradicting $\lambda\in\sigma(L)$. This proves $\sigma(L)\subset\Lambda_{\varepsilon_{K}}(L_{K})$.
For the multiplicity statement, the definition of $U$ implies $\sup_{z\in\partial U}\mathcal{R}(z,L_{K})<1/\varepsilon_{K}$, hence $\partial U\subset\rho(L)$ by Corollary 2.3. Consider $L_{t}:=L_{K}+t(L-L_{K})$, $t\in[0,1]$. For $z\in\partial U$,
$\|(L_{t}-L_{K})(zI-L_{K})^{-1}\|\leq t\,\varepsilon_{K}\,\mathcal{R}(z,L_{K})<1,$
so $\partial U\subset\rho(L_{t})$ for all $t$. Therefore $P_{L_{t}}(U)$ varies continuously with $t$ in operator norm, and $\dim\operatorname{Ran}P_{L_{t}}(U)$ is constant. Evaluating at $t=0$ and $t=1$ gives the claim. ∎
2.2 Resolvent perturbation bound and spectral exclusions
Definition 2.5.
Let $A\in\mathcal{L}(\mathcal{B})$ and let $\Gamma$ be a piecewise $C^{1}$ closed curve in $\mathbb{C}$. We say that $\Gamma$ is a resolvent curve for $A$ if $\Gamma\subset\rho(A)$. If $\Gamma$ bounds a domain $U$ and $\Gamma\subset\rho(A)$, we also call $U$ a spectral window for $A$.
Proposition 2.6.
Assume $\|L-L_{K}\|\leq\varepsilon_{K}$ and let $\Gamma\subset\rho(L_{K})$ be a closed curve such that
$\sup_{z\in\Gamma}\ \varepsilon_{K}\,\mathcal{R}(z,L_{K})\ <\ 1.$
Then $\Gamma\subset\rho(L)$ as well. In particular, if $\Gamma=\partial U$ for a bounded domain $U$, then $\sigma(L)$ cannot intersect $\Gamma$ and the Riesz projector $P_{L}(U)$ is well defined.
Proof.
For each $z\in\Gamma$ we have
$\|(L-L_{K})R_{L_{K}}(z)\|\leq\varepsilon_{K}\,\mathcal{R}(z,L_{K})<1,$
so $z\in\rho(L)$ by Corollary 2.3. ∎
2.3 Riesz projectors, multiplicity, and invariant subspaces
Definition 2.7.
Let $A\in\mathcal{L}(\mathcal{B})$ and let $U\subset\mathbb{C}$ be a bounded domain such that $\Gamma:=\partial U\subset\rho(A)$ is a piecewise $C^{1}$ Jordan curve. The Riesz projector onto the spectral subset $\sigma(A)\cap U$ is
$P_{A}(U):=\frac{1}{2\pi i}\int_{\Gamma}(zI-A)^{-1}\,dz.$
Its range $\mathcal{E}_{A}(U):=\operatorname{Ran}P_{A}(U)$ is $A$-invariant. The total algebraic multiplicity of eigenvalues of $A$ in $U$ is $\dim\mathcal{E}_{A}(U)$ (finite in the compact setting when $\textrm{dist}(U,0)>0$).
Assumption 2.8.
Let $L,L_{K}\in\mathcal{L}(\mathcal{B})$ satisfy $\|L-L_{K}\|\leq\varepsilon_{K}$. Let $U\subset\mathbb{C}$ be a bounded domain with piecewise $C^{1}$ Jordan boundary $\Gamma:=\partial U$ such that $\Gamma\subset\rho(L_{K})$ and
$\eta:=\sup_{z\in\Gamma}\ \|(L-L_{K})R_{L_{K}}(z)\|\leq\varepsilon_{K}\,\sup_{z\in\Gamma}\mathcal{R}(z,L_{K})<1.$
Lemma 2.9.
Assume Assumption 2.8. Then $\Gamma\subset\rho(L)$ and, writing
$P:=P_{L}(U),\qquad P_{K}:=P_{L_{K}}(U),\qquad M:=\sup_{z\in\Gamma}\mathcal{R}(z,L_{K}),$
we have
$\|P-P_{K}\|\leq\frac{\ell(\Gamma)}{2\pi}\cdot\frac{\varepsilon_{K}\,M^{2}}{1-\eta},$
where $\ell(\Gamma)$ denotes the length of $\Gamma$.
Proof.
Fix $z\in\Gamma$. Since $z\in\rho(L_{K})$ and $\|(L-L_{K})R_{L_{K}}(z)\|\leq\eta<1$, the Neumann factorization
$zI-L=\bigl(I-(L-L_{K})R_{L_{K}}(z)\bigr)\,(zI-L_{K})$
shows $z\in\rho(L)$ and
$R_{L}(z)=R_{L_{K}}(z)\bigl(I-(L-L_{K})R_{L_{K}}(z)\bigr)^{-1}.$
Hence
$\|R_{L}(z)\|\leq\frac{\|R_{L_{K}}(z)\|}{1-\|(L-L_{K})R_{L_{K}}(z)\|}\leq\frac{\mathcal{R}(z,L_{K})}{1-\eta}.$
Using the resolvent identity
$R_{L}(z)-R_{L_{K}}(z)=R_{L}(z)\,(L-L_{K})\,R_{L_{K}}(z),$
we obtain
$\|R_{L}(z)-R_{L_{K}}(z)\|\leq\|R_{L}(z)\|\,\|L-L_{K}\|\,\|R_{L_{K}}(z)\|\leq\frac{\varepsilon_{K}\,\mathcal{R}(z,L_{K})^{2}}{1-\eta}.$
Finally,
$\|P-P_{K}\|=\left\|\frac{1}{2\pi i}\int_{\Gamma}\bigl(R_{L}(z)-R_{L_{K}}(z)\bigr)\,dz\right\|\leq\frac{1}{2\pi}\int_{\Gamma}\|R_{L}(z)-R_{L_{K}}(z)\|\,|dz|$
and bounding $\mathcal{R}(z,L_{K})\leq M$ along $\Gamma$ yields the claim. ∎
Remark 2.10.
If $\|P-P_{K}\|<1$, then $P$ restricts to an isomorphism $\operatorname{Ran}P_{K}\to\operatorname{Ran}P$; in particular, $\dim\operatorname{Ran}P=\dim\operatorname{Ran}P_{K}$.
Lemma 2.11.
Assume Assumption 2.8 and suppose
$\operatorname{rank}P=\operatorname{rank}P_{K}=1,\qquad\vartheta:=\|P-P_{K}\|<1.$
Let $v_{K}\in\operatorname{Ran}P_{K}$ with $\|v_{K}\|=1$. Then
$\|(I-P)v_{K}\|\leq\vartheta,\qquad\|Pv_{K}\|\geq 1-\vartheta.$
In particular, the normalized vector
$v:=\frac{Pv_{K}}{\|Pv_{K}\|}\in\operatorname{Ran}P$
is well-defined and satisfies
$\|v-v_{K}\|\leq\frac{2\vartheta}{1-\vartheta}.$
Proof.
Since $v_{K}\in\operatorname{Ran}P_{K}$ we have $P_{K}v_{K}=v_{K}$, hence
$(I-P)v_{K}=(P_{K}-P)v_{K},\qquad\Rightarrow\qquad\|(I-P)v_{K}\|\leq\|P-P_{K}\|\,\|v_{K}\|=\vartheta.$
Thus
$\|Pv_{K}\|\geq\|v_{K}\|-\|(I-P)v_{K}\|\geq 1-\vartheta,$
so $v$ is well-defined. Next,
$\|v-v_{K}\|=\left\|\frac{Pv_{K}}{\|Pv_{K}\|}-v_{K}\right\|\leq\left\|\frac{Pv_{K}}{\|Pv_{K}\|}-Pv_{K}\right\|+\|Pv_{K}-v_{K}\|.$
The second term is $\|Pv_{K}-v_{K}\|=\|(P-P_{K})v_{K}\|\leq\vartheta$. For the first term, $\bigl\|\frac{Pv_{K}}{\|Pv_{K}\|}-Pv_{K}\bigr\|=\bigl|1-\|Pv_{K}\|\bigr|\,\|Pv_{K}\|^{-1}\,\|Pv_{K}\|=\bigl|1-\|Pv_{K}\|\bigr|\leq\|v_{K}-Pv_{K}\|=\|(I-P)v_{K}\|\leq\vartheta$, and dividing by $\|Pv_{K}\|\geq 1-\vartheta$ yields
$\|v-v_{K}\|\leq\frac{\vartheta+\vartheta}{1-\vartheta}=\frac{2\vartheta}{1-\vartheta}.$
Lemma 2.12.
Assume the setting of Lemma 2.11 and moreover $\|L_{K}\|\leq C$. Let $\lambda$ be the (unique) eigenvalue of $L$ in $U$ and $\lambda_{K}$ the (unique) eigenvalue of $L_{K}$ in $U$. Then
$|\lambda-\lambda_{K}|\ \leq\ \frac{\varepsilon_{K}(1+\vartheta)+2C\,\vartheta}{1-\vartheta}.$
Proof.
Since $\operatorname{rank}P=\operatorname{rank}P_{K}=1$, both $\operatorname{Ran}P$ and $\operatorname{Ran}P_{K}$ are one-dimensional invariant subspaces, and $L|_{\operatorname{Ran}P}$ (resp. $L_{K}|_{\operatorname{Ran}P_{K}}$) acts as multiplication by $\lambda$ (resp. $\lambda_{K}$).
Define the linear map
$\Phi:=P|_{\operatorname{Ran}P_{K}}:\operatorname{Ran}P_{K}\to\operatorname{Ran}P.$
For $x\in\operatorname{Ran}P_{K}$, we have $x=P_{K}x$, hence
$\|\Phi x-x\|=\|(P-P_{K})x\|\leq\vartheta\|x\|,\qquad\Rightarrow\qquad\|\Phi x\|\geq(1-\vartheta)\|x\|.$
Therefore $\Phi$ is invertible and $\|\Phi^{-1}\|\leq(1-\vartheta)^{-1}$. Also, on $\operatorname{Ran}P_{K}$,
$\|\Phi x\|=\|Px\|\leq\|x\|+\|(P-P_{K})x\|\leq(1+\vartheta)\|x\|,\qquad\Rightarrow\qquad\|\Phi\|\leq 1+\vartheta.$
Now observe that for $x\in\operatorname{Ran}P_{K}$,
$\Phi^{-1}L\Phi\,x\in\operatorname{Ran}P_{K},\qquad\text{and}\qquad\Phi^{-1}L\Phi\,x=\lambda\,x,$
because $\Phi x\in\operatorname{Ran}P$ and $L$ acts as $\lambda$ on $\operatorname{Ran}P$, while $\Phi^{-1}$ maps $\operatorname{Ran}P$ back to $\operatorname{Ran}P_{K}$. Hence, as operators on the one-dimensional space $\operatorname{Ran}P_{K}$,
$|\lambda-\lambda_{K}|=\bigl\|\Phi^{-1}L\Phi-L_{K}|_{\operatorname{Ran}P_{K}}\bigr\|.$
Insert $L=L_{K}+(L-L_{K})$ and add and subtract $\Phi^{-1}L_{K}\Phi$:
$\|\Phi^{-1}L\Phi-L_{K}\|\leq\|\Phi^{-1}(L-L_{K})\Phi\|+\|\Phi^{-1}L_{K}\Phi-L_{K}\|.$
The first term satisfies
$\|\Phi^{-1}(L-L_{K})\Phi\|\leq\|\Phi^{-1}\|\,\|L-L_{K}\|\,\|\Phi\|\leq\frac{\varepsilon_{K}(1+\vartheta)}{1-\vartheta}.$
For the second term, note that on $\operatorname{Ran}P_{K}$ we have $\Phi=P$ and $P_{K}=\mathrm{Id}$, hence
$\Phi^{-1}L_{K}\Phi-L_{K}=\Phi^{-1}\bigl(L_{K}P-PL_{K}\bigr).$
Since $\Phi^{-1}$ is bounded by $(1-\vartheta)^{-1}$, it remains to bound the commutator:
$\|L_{K}P-PL_{K}\|\leq\|L_{K}(P-P_{K})\|+\|(P-P_{K})L_{K}\|\leq 2\|L_{K}\|\,\|P-P_{K}\|\leq 2C\,\vartheta.$
Putting the two bounds together gives
$|\lambda-\lambda_{K}|\leq\frac{\varepsilon_{K}(1+\vartheta)+2C\,\vartheta}{1-\vartheta}.$
Remark 2.13.
A certified bound on $\|P-P_{K}\|$ yields: (i) a certified invariant subspace enclosure $\operatorname{Ran}P_{K}\approx\operatorname{Ran}P$ (always), (ii) in the rank-one (simple eigenvalue) case, an explicit bound on the eigenvector direction (Lemma 2.11), and (iii) combined with $\|L-L_{K}\|$ (and a crude bound on $\|L_{K}\|$), an explicit eigenvalue error bound (Lemma 2.12). For higher multiplicity, $P$ certifies the whole invariant subspace; individual eigenvectors require additional structure (e.g. normality, a chosen basis, or a spectral gap within the cluster).
2.4 Coarse–fine approach
Under Assumption 2.1, for all $K^{\prime}\geq K$,
$\|L_{K^{\prime}}-L_{K}\|\ \leq\ \varepsilon_{K},\qquad\text{and}\qquad\|L-L_{K^{\prime}}\|\ \leq\ \varepsilon_{K^{\prime}}\ \leq\ \varepsilon_{K}.$ (1)
Thus a resolvent estimate at a coarse level $K$ propagates to any finer $K^{\prime}\geq K$; the DFLY analogue requires additional bookkeeping (Appendix A).
Fix a region $U\subset\mathbb{C}$ and assume that $\partial U\subset\rho(L_{K})$ and that
$\sup_{z\in\partial U}\ \varepsilon_{K}\,\mathcal{R}(z,L_{K})\ <\ 1.$ (2)
Then, by applying Corollary 2.3 with $A=L_{K^{\prime}}$ and $B=L_{K}$ and using (1), we obtain $\partial U\subset\rho(L_{K^{\prime}})$ and the explicit bound
$\mathcal{R}(z,L_{K^{\prime}})\ \leq\ \frac{\mathcal{R}(z,L_{K})}{1-\varepsilon_{K}\,\mathcal{R}(z,L_{K})}\qquad\text{for all }z\in\partial U.$ (3)
In particular, the same coarse-level contour $\partial U$ that isolates (say) a simple eigenvalue of $L_{K}$ also isolates the corresponding eigenvalue of $L_{K^{\prime}}$ for every $K^{\prime}\geq K$, with no need to recompute resolvents from scratch at the fine level.
This yields a coarse-fine workflow:
  1. 1.
    At a modest $K$, compute $\mathcal{R}(z,L_{K})$ and use Proposition 2.4 to produce enclosures for $\sigma(L)$ and certify spectral separation.
  2. 2.
    Increase to $K^{\prime}\gg K$. Use (3) to control $\mathcal{R}(z,L_{K^{\prime}})$ on the same contour $\partial U$, then apply Lemma 2.9 with the smaller error $\varepsilon_{K^{\prime}}$ to obtain tighter projector bounds.
Remark 2.14.
The two levels play different roles. At the coarse level $K$, the resolvent bounds need only be good enough to certify that the contour $\partial U$ lies in $\rho(L)$, i.e., to establish spectral isolation. The resulting projector bound $\|P-P_{K}\|$ may be far too large for useful eigenvector or eigenvalue enclosures. At the fine level $K^{\prime}$, the same contour is reused and the projector bound improves by a factor of $\varepsilon_{K^{\prime}}/\varepsilon_{K}$, without recomputing any resolvent. In particular, the coarse matrix can be small (and hence cheap to certify), while the fine matrix is large and delivers precision.
The following theorem shows that the perturbation condition is not merely a checkable sufficient condition: it is eventually satisfied, so the certification pipeline is guaranteed to succeed at some finite truncation level.
Theorem 2.15.
Under Assumption 2.1, let $U\subset\mathbb{C}$ be a bounded domain with piecewise $C^{1}$ boundary $\Gamma:=\partial U$ such that $0\notin\overline{U}$, $\Gamma\subset\rho(L)$, and $\sigma(L)\cap\Gamma=\emptyset$. Then there exists $K_{0}$ such that for all $K\geq K_{0}$:
  1. 1.
    $\Gamma\subset\rho(L_{K})$ and the perturbation condition $\sup_{z\in\Gamma}\varepsilon_{K}\,\mathcal{R}(z,L_{K})<1$ holds;
  2. 2.
    $\dim\operatorname{Ran}P_{L_{K}}(U)=\dim\operatorname{Ran}P_{L}(U)$;
  3. 3.
    $\|P_{L_{K}}(U)-P_{L}(U)\|\to 0$ as $K\to\infty$.
In particular, every eigenvalue, eigenvector, and Riesz projector of $L$ in $U$ can be approximated to arbitrary precision by the finite-rank data of $L_{K}$.
Proof.
Since $\Gamma$ is compact and $\Gamma\subset\rho(L)$, we have $M:=\sup_{z\in\Gamma}\mathcal{R}(z,L)<\infty$. Because $\|L-L_{K}\|\leq\varepsilon_{K}\to 0$, the second resolvent identity gives $\mathcal{R}(z,L_{K})\to\mathcal{R}(z,L)$ uniformly on $\Gamma$. In particular, $\sup_{z\in\Gamma}\varepsilon_{K}\,\mathcal{R}(z,L_{K})\to 0$, so the perturbation condition holds for $K\geq K_{0}$. Claims (ii) and (iii) then follow from Proposition 2.4 and Lemma 2.9. ∎
3 Certified resolvent bounds for discretizations (validated linear algebra)
Let $A_{K}\in\mathbb{C}^{N\times N}$ represent $L_{K}|_{E_{K}}$ in a chosen basis of $E_{K}=\operatorname{Ran}\pi_{K}$. Throughout this section $\|\cdot\|$ denotes the Euclidean operator norm $\|\cdot\|_{2}$.
3.1 Schur/SVD-based resolvent bounds with explicit defects
A certified Schur build and its defects
Let $A\in\mathbb{C}^{N\times N}$ denote $A_{K}$ for simplicity. Compute a complex Schur decomposition (numerically):
$AQ=QT+R,$ (4)
where $Q\in\mathbb{C}^{N\times N}$ is approximately unitary, $T\in\mathbb{C}^{N\times N}$ is upper triangular, and $R$ is the Schur residual. Define the defects
$\delta:=\|I-Q^{\ast}Q\|,\qquad r_{\mathrm{sch}}:=\|AQ-QT\|=\|R\|,\qquad C_{A}:=\|A\|.$ (5)
Define the proxy matrix
$S_{0}:=QTQ^{\ast},\qquad E:=A-S_{0}.$ (6)
Lemma 3.1 (cf. [12]).
With $S_{0}$ and $E$ defined by (6), we have
$\|E\|\ \leq\ r_{\mathrm{sch}}\sqrt{1+\delta}\ +\ C_{A}\,\delta.$ (7)
Moreover, if $\delta<1$ then $Q$ is invertible and
$\|Q\|\leq\sqrt{1+\delta},\qquad\|Q^{-1}\|\leq(1-\delta)^{-1/2},\qquad\kappa(Q):=\|Q\|\,\|Q^{-1}\|\leq\sqrt{\frac{1+\delta}{1-\delta}}.$ (8)
Proof.
Start from $E=A-QTQ^{\ast}$ and add/subtract $AQQ^{\ast}$:
$E=(AQ-QT)Q^{\ast}\ +\ A(I-QQ^{\ast}).$
Taking norms and using $\|Q\|\leq\sqrt{1+\delta}$ (from $(1-\delta)I\preceq Q^{\ast}Q\preceq(1+\delta)I$) and $\|I-QQ^{\ast}\|=\|I-Q^{\ast}Q\|=\delta$ gives (7). The bounds (8) follow from the same Loewner inequality. ∎
Resolvent bounds via a Neumann argument
Lemma 3.2.
Fix $z\in\mathbb{C}$. If
$\alpha(z):=\|(zI-S_{0})^{-1}\|\,\|E\|<1,$
then $z\in\rho(A)$ and
$\|(zI-A)^{-1}\|\ \leq\ \frac{\|(zI-S_{0})^{-1}\|}{1-\alpha(z)}.$ (9)
Proof.
Factor $zI-A=(I-(zI-S_{0})^{-1}E)\,(zI-S_{0})$ and invert via Neumann series. ∎
Lemma 3.3.
Assume $\delta=\|I-Q^{\ast}Q\|<1$ and let $\kappa(Q)$ be as in (8). Then for every $z\in\mathbb{C}$,
$\|(zI-S_{0})^{-1}\|\ \leq\ \kappa(Q)^{2}\ \|(zI-T)^{-1}\|.$ (10)
Proof.
Let $Q=UH$ be the polar decomposition. Then $S_{0}=U(HTH)U^{\ast}$, so $(zI-S_{0})^{-1}=U\,H^{-1}(zI-T)^{-1}H^{-1}\,U^{\ast}$ and $\|(zI-S_{0})^{-1}\|\leq\|H^{-1}\|^{2}\,\|(zI-T)^{-1}\|\leq\kappa(Q)^{2}\,\|(zI-T)^{-1}\|$. ∎
Theorem 3.4.
Let $A\in\mathbb{C}^{N\times N}$ and suppose (4) holds with defects $\delta,r_{\mathrm{sch}},C_{A}$ as in (5). Assume $\delta<1$ and define $S_{0},E$ by (6). Then
$\|E\|\ \leq\ r_{\mathrm{sch}}\sqrt{1+\delta}+C_{A}\,\delta,\qquad\|(zI-S_{0})^{-1}\|\ \leq\ \kappa(Q)^{2}\,\|(zI-T)^{-1}\|.$
Consequently, if
$\beta(z):=\kappa(Q)^{2}\,\|(zI-T)^{-1}\|\,\|E\|<1,$
then $z\in\rho(A)$ and
$\|(zI-A)^{-1}\|\ \leq\ \frac{\kappa(Q)^{2}\,\|(zI-T)^{-1}\|}{1-\beta(z)}.$ (11)
Proof.
Combine Lemma 3.1, Lemma 3.3, and Lemma 3.2. ∎
3.2 Triangular resolvent bounds: SVD and block splitting
Since $\|(zI-T)^{-1}\|=1/\sigma_{\min}(zI-T)$, it suffices to produce certified lower bounds for $\sigma_{\min}(zI-T)$. The following result from validated linear algebra provides such bounds.
Theorem 3.5 ([50, Thm. 3.1]).
Let $A,U,V\in\mathbb{K}^{n\times n}$ with $\|I-U^{H}U\|\leq\alpha<1$ and $\|I-V^{H}V\|\leq\beta<1$. Write $U^{H}AV=D+E$ with $D$ diagonal. Then there exists a permutation $\nu$ such that
$\frac{|D_{ii}|-\|E\|}{\sqrt{(1+\alpha)(1+\beta)}}\;\leq\;\sigma_{\nu(i)}(A)\;\leq\;\frac{|D_{ii}|+\|E\|}{\sqrt{(1-\alpha)(1-\beta)}}.$
In practice we apply Theorem 3.5 to A=zITA=zI-T; see [46] for extensions and [12] for the implementation.
Block splitting near a target eigenvalue
When $z$ is close to a cluster of diagonal entries of $T$, computing $\sigma_{\min}(zI-T)$ at high precision for the full $N\times N$ matrix can be wasteful. Instead, we isolate the difficult part in a small leading block.
Let $T$ be Schur-ordered so that the $k$ diagonal entries closest to a target $\lambda_{j}$ appear first, and partition
$T=\begin{pmatrix}T_{11}&T_{12}\\ 0&T_{22}\end{pmatrix}$.
Lemma 3.6.
For every $z\in\mathbb{C}$,
$\sigma_{\min}(zI-T)\ \geq\ \min\bigl(\sigma_{\min}(zI-T_{11}),\ \sigma_{\min}(zI-T_{22})\bigr)\ -\ \|T_{12}\|.$
Proof.
Write zIT=M(z)+EzI-T=M(z)+E with
$M(z):=\begin{pmatrix}zI-T_{11}&0\\ 0&zI-T_{22}\end{pmatrix},\qquad E:=\begin{pmatrix}0&-T_{12}\\ 0&0\end{pmatrix}$.
Then $\sigma_{\min}(M(z))=\min(\sigma_{\min}(zI-T_{11}),\sigma_{\min}(zI-T_{22}))$ and $\|E\|=\|T_{12}\|$. Apply Weyl: $\sigma_{\min}(M(z)+E)\geq\sigma_{\min}(M(z))-\|E\|$. ∎
Sampling a contour and Lipschitz propagation
Lemma 3.7.
For any fixed matrix $A$ and any $z,w\in\mathbb{C}$,
$\bigl|\sigma_{\min}(zI-A)-\sigma_{\min}(wI-A)\bigr|\leq|z-w|.$
Proof.
Since (zIA)(wIA)=(zw)I(zI-A)-(wI-A)=(z-w)I and $\|(z-w)I\|=|z-w|$, Weyl’s inequality gives $|\sigma_{\min}(zI-A)-\sigma_{\min}(wI-A)|\leq\|(z-w)I\|=|z-w|$. ∎
For a circle $\Gamma=\{\,\lambda+\rho e^{it}\,\}$ sampled at $m$ equally spaced points, every point of the circle lies within chord distance
$\Delta_{m}:=2\rho\sin\!\Bigl(\frac{\pi}{2m}\Bigr)$
of some sample point.
Proposition 3.8.
Let $T$ be upper triangular and let $\Gamma=\{z:\ |z-\lambda|=\rho\}$. Fix sample points $z_{\ell}=\lambda+\rho e^{2\pi i\ell/m}$, $\ell=0,\dots,m-1$, and let
$s_{\ell}\ \leq\ \sigma_{\min}(z_{\ell}I-T)\qquad(\ell=0,\dots,m-1)$
be certified lower bounds. Set
$s_{*}:=\min_{0\leq\ell<m}s_{\ell}\ -\ \Delta_{m},\qquad\Delta_{m}:=2\rho\sin\!\Bigl(\frac{\pi}{2m}\Bigr).$
If $s_{*}>0$, then $\Gamma\subset\rho(T)$ and
$\sup_{z\in\Gamma}\ \|(zI-T)^{-1}\|\ \leq\ \frac{1}{s_{*}}.$
Proof.
For any $z\in\Gamma$ choose $\ell$ with $|z-z_{\ell}|\leq\Delta_{m}$ and apply Lemma 3.7:
$\sigma_{\min}(zI-T)\geq\sigma_{\min}(z_{\ell}I-T)-|z-z_{\ell}|\geq s_{\ell}-\Delta_{m}\geq s_{*}.$
Thus $\sigma_{\min}(zI-T)\geq s_{*}>0$ for all $z\in\Gamma$, hence all zITzI-T are invertible and $\|(zI-T)^{-1}\|=1/\sigma_{\min}(zI-T)\leq 1/s_{*}$. ∎
Remark 3.9.
In [12] the equispaced sampling of Proposition 3.8 is replaced by an adaptive method that refines the mesh where $\sigma_{\min}(z_{\ell}I-T)$ is small, reducing the number of certified SVD evaluations while maintaining the same Lipschitz-propagation guarantee.
Remark 3.10.
In Proposition 3.8, the bounds $s_{\ell}$ can be produced using Lemma 3.6: compute a certified $c_{12}\geq\|T_{12}\|$ once, then at each sample point bound $\sigma_{\min}(z_{\ell}I-T_{11})$ (high precision, small $k$) and $\sigma_{\min}(z_{\ell}I-T_{22})$ (fast, since $T_{22}$ is well separated), and combine
$s_{\ell}:=\min\bigl(s^{(11)}_{\ell},\ s^{(22)}_{\ell}\bigr)-c_{12}.$
3.3 Contour evaluation and certified suprema
We summarize the certification workflow for producing $M_{\Gamma}(A):=\sup_{z\in\Gamma}\|(zI-A)^{-1}\|$ on a contour $\Gamma=\partial U$.
  1. 1.
    Compute $(Q,T)$ and the defects $\delta,r_{\mathrm{sch}},C_{A}$ in (5). Form the bound on $\|E\|$ from Lemma 3.1 and the conditioning factor $\kappa(Q)$.
  2. 2.
    Choose a sampling of $\Gamma$ (e.g. circles around selected diagonal entries of $T$). At sample points $z_{\ell}$, produce certified lower bounds for $\sigma_{\min}(z_{\ell}I-T)$ (either directly by SVD, or by block splitting as in Remark 3.10). Use Proposition 3.8 (or the obvious analogue for a general piecewise $C^{1}$ contour) to obtain a bound
    $M_{\Gamma}(T)\ :=\ \sup_{z\in\Gamma}\|(zI-T)^{-1}\|.$
  3. 3.
    Define
    $\beta_{\Gamma}:=\kappa(Q)^{2}\,M_{\Gamma}(T)\,\|E\|.$
    If $\beta_{\Gamma}<1$, then Theorem 3.4 gives $\Gamma\subset\rho(A)$ and
    $M_{\Gamma}(A)\ \leq\ \frac{\kappa(Q)^{2}\,M_{\Gamma}(T)}{1-\beta_{\Gamma}}.$
If $\varepsilon_{K}\,M_{\Gamma}(A)<1$, then $\Gamma$ is a certified exclusion curve for the original operator $L$.
3.4 Projector error decomposition
Fix a bounded domain $U\subset\mathbb{C}$ whose boundary $\Gamma:=\partial U$ is a piecewise $C^{1}$ Jordan curve. For any matrix $M$, write the Riesz projector
$P_{M}(U):=\frac{1}{2\pi i}\int_{\Gamma}(zI-M)^{-1}\,dz.$
Let $A\in\mathbb{C}^{N\times N}$ and suppose we have a (numerical) Schur pair $(Q,T)$ with
$AQ=QT+R_{\mathrm{sch}},\qquad r_{\mathrm{sch}}:=\|R_{\mathrm{sch}}\|_{2},$
where $Q$ is unitary (or already certified as nearly unitary). Define the Schur-basis representation
$\widehat{A}:=Q^{\ast}AQ=T+E_{\mathrm{sch}},\qquad E_{\mathrm{sch}}:=Q^{\ast}R_{\mathrm{sch}},\qquad\|E_{\mathrm{sch}}\|_{2}=r_{\mathrm{sch}}.$
Next, suppose a reordering routine produces $\widehat{U}$ and $\tilde{T}$ such that
$T\,\widehat{U}=\widehat{U}\,\tilde{T}+R_{\mathrm{ord}},\qquad\delta_{\mathrm{ord}}:=\|R_{\mathrm{ord}}\|_{2},$
and let the orthogonality defect be
$\Delta_{U}:=I-\widehat{U}^{\ast}\widehat{U},\qquad\delta_{U}:=\|\Delta_{U}\|_{2}.$
Assume $\delta_{U}<1$ so that $\widehat{U}$ is invertible and $\|\widehat{U}^{-1}\|_{2}\leq(1-\delta_{U})^{-1/2}$.
Since $\tilde{T}$ is expressed in the reordered basis, we compare projectors in the original Schur coordinates via
$\tilde{P}_{T}(U):=\widehat{U}\,P_{\tilde{T}}(U)\,\widehat{U}^{-1}.$
Lemma 3.11.
With the objects above,
$\|P_{\widehat{A}}(U)-\tilde{P}_{T}(U)\|\leq\|P_{\widehat{A}}(U)-P_{T}(U)\|\;+\;\|P_{T}(U)-\widehat{U}P_{\tilde{T}}(U)\widehat{U}^{-1}\|.$
Proof.
Insert and subtract $P_{T}(U)$ and use the triangle inequality. ∎
The first term in the triangle inequality is controlled by the Schur residual. Let
$M_{T}:=\sup_{z\in\Gamma}\|(zI-T)^{-1}\|_{2}.$
If
$\eta_{\mathrm{sch}}:=\|E_{\mathrm{sch}}\|_{2}\,M_{T}=r_{\mathrm{sch}}\,M_{T}<1,$
[…]
$\sup_{z\in\Gamma}\|(zI-\widehat{A})^{-1}-(zI-T)^{-1}\|_{2}\leq\frac{r_{\mathrm{sch}}\,M_{T}^{2}}{1-\eta_{\mathrm{sch}}}.$
Consequently,
$\|P_{\widehat{A}}(U)-P_{T}(U)\|\leq\frac{\ell(\Gamma)}{2\pi}\cdot\frac{r_{\mathrm{sch}}\,M_{T}^{2}}{1-\eta_{\mathrm{sch}}},$ (12)
where $\ell(\Gamma)$ denotes the length of $\Gamma$.
The second term is controlled by the reordering defect. From $T\widehat{U}=\widehat{U}\tilde{T}+R_{\mathrm{ord}}$ we have
$T=\widehat{U}\tilde{T}\,\widehat{U}^{-1}+E_{\mathrm{ord}},\qquad E_{\mathrm{ord}}:=R_{\mathrm{ord}}\,\widehat{U}^{-1},\qquad\|E_{\mathrm{ord}}\|_{2}\leq\delta_{\mathrm{ord}}\,(1-\delta_{U})^{-1/2}.$
If
$\eta_{\mathrm{ord}}:=\|E_{\mathrm{ord}}\|_{2}\,M_{T}<1,$
then $\Gamma$ also lies in the resolvent set of $\widehat{U}\tilde{T}\widehat{U}^{-1}$ and
$\sup_{z\in\Gamma}\|(zI-T)^{-1}-(zI-\widehat{U}\tilde{T}\widehat{U}^{-1})^{-1}\|_{2}\leq\frac{\|E_{\mathrm{ord}}\|_{2}\,M_{T}^{2}}{1-\eta_{\mathrm{ord}}}.$
Using similarity invariance of the Riesz projector in exact arithmetic, $P_{\widehat{U}\tilde{T}\widehat{U}^{-1}}(U)=\widehat{U}P_{\tilde{T}}(U)\widehat{U}^{-1}$, we obtain
$\|P_{T}(U)-\widehat{U}P_{\tilde{T}}(U)\widehat{U}^{-1}\|\leq\frac{\ell(\Gamma)}{2\pi}\cdot\frac{\|E_{\mathrm{ord}}\|_{2}\,M_{T}^{2}}{1-\eta_{\mathrm{ord}}}.$ (13)
Under $\eta_{\mathrm{sch}}<1$ and $\eta_{\mathrm{ord}}<1$, Lemma 3.11 together with (12)–(13) yields an explicit certified estimate for $\|P_{\widehat{A}}(U)-\tilde{P}_{T}(U)\|$. Finally, since $P_{A}(U)=Q\,P_{\widehat{A}}(U)\,Q^{\ast}$ and $Q$ is unitary, this also bounds the projector error in the original coordinates of $A$.
4 Hardy-space framework for the Gauss–Kuzmin–Wirsing operator
4.1 Hardy spaces and branch geometry
For $r>0$ let
$D_{r}:=\{\,w\in\mathbb{C}:\ |w-1|<r\,\}$
and let $H^{2}(D_{r})$ be the Hardy space on $D_{r}$, realized as holomorphic functions on $D_{r}$ with square summable Taylor coefficients at $1$.
Definition 4.1.
For $r>0$, define
$H^{2}(D_{r}):=\left\{\,f(w)=\sum_{k\geq 0}c_{k}(w-1)^{k}\ \text{holomorphic on }D_{r}:\ \|f\|_{H^{2}(D_{r})}^{2}:=\sum_{k\geq 0}|c_{k}|^{2}r^{2k}<\infty\right\}.$
With the inner product $\langle f,g\rangle_{H^{2}(D_{r})}:=\sum_{k\geq 0}c_{k}\overline{d_{k}}\,r^{2k}$, $H^{2}(D_{r})$ is a Hilbert space. If $0<r<R$, restriction defines a continuous embedding
$J_{R\to r}:H^{2}(D_{R})\hookrightarrow H^{2}(D_{r}),\qquad(J_{R\to r}f)(w)=f(w)\big|_{D_{r}},$
with operator norm $\|J_{R\to r}\|=1$.
The Gauss inverse branches are
$\tau_{n}(w)=\frac{1}{w+n},\qquad n\geq 1.$
They map discs $D_{3/2}$ strictly inside $D_{1}$, which is the key geometric fact behind the strong-weak smoothing that later yields compactness and explicit approximation bounds.
Lemma 4.2.
For every $n\geq 1$ one has
$\tau_{n}(D_{3/2})\subset D_{1}.$
Proof.
If $w\in D_{3/2}$, then $\Re w>-\tfrac{1}{2}$. Compute
$|\tau_{n}(w)-1|=\left|\frac{1}{w+n}-1\right|=\frac{|w+n-1|}{|w+n|}.$
Moreover
$|w+n|^{2}-|w+n-1|^{2}=(w+n)(\overline{w}+n)-(w+n-1)(\overline{w}+n-1)=2\Re w+(2n-1).$
Since $\Re w>-\tfrac{1}{2}$, we get $2\Re w+(2n-1)>2n-2\geq 0$. Hence |w+n|>|w+n1||w+n|>|w+n-1| and therefore $|\tau_{n}(w)-1|<1$, i.e. $\tau_{n}(w)\in D_{1}$. ∎
Lemma 4.3.
For every $n\geq 1$ and every $w\in D_{3/2}$,
$|(w+n)^{-2}|\ \leq\ (n-\tfrac{1}{2})^{-2}.$
Proof.
If $w\in D_{3/2}$ then $\Re w>-\tfrac{1}{2}$ and hence $|w+n|>n-\tfrac{1}{2}$, which implies the claim. ∎
4.2 Boundedness and compactness
Define the (un-normalized) Gauss–Kuzmin–Wirsing transfer operator by
$(\mathcal{L}f)(w):=\sum_{n\geq 1}\frac{1}{(w+n)^{2}}\,f\!\left(\frac{1}{w+n}\right)=\sum_{n\geq 1}\psi_{n}(w)\,(f\circ\tau_{n})(w),\qquad\psi_{n}(w):=(w+n)^{-2}.$ (14)
The geometric inclusion of Lemma 4.2 ensures that if $f$ is holomorphic on $D_{1}$, then $f\circ\tau_{n}$ is holomorphic on $D_{3/2}$, so each summand in (14) is holomorphic on $D_{3/2}$.
Proposition 4.4.
The operator $\mathcal{L}$ defined by (14) extends to a bounded operator
$\mathcal{L}:H^{2}(D_{1})\longrightarrow H^{2}(D_{3/2}),$
and
$\|\mathcal{L}\|_{H^{2}(D_{1})\to H^{2}(D_{3/2})}\ \leq\ \sum_{n\geq 1}\frac{\|C_{\tau_{n}}\|_{H^{2}(D_{1})\to H^{2}(D_{3/2})}}{(n-\tfrac{1}{2})^{2}}=:C_{2},$
where $C_{\tau_{n}}:f\mapsto f\circ\tau_{n}$ is the composition operator. In particular, using the estimate $\|C_{\tau_{n}}\|\leq\sqrt{2n+1}$ (proved below),
$\|\mathcal{L}\|_{H^{2}(D_{1})\to H^{2}(D_{3/2})}\ \leq\ \sum_{n\geq 1}\frac{\sqrt{2n+1}}{(n-\tfrac{1}{2})^{2}}\ \leq\ \sqrt{6}\,\bigl(2^{3/2}-1\bigr)\,\zeta\!\left(\tfrac{3}{2}\right)=:C_{2}.$
Proof.
Fix $f\in H^{2}(D_{1})$. For each $n\geq 1$, the map $w\mapsto(f\circ\tau_{n})(w)$ is holomorphic on $D_{3/2}$ by Lemma 4.2, and multiplication by the bounded holomorphic weight $\psi_{n}$ preserves holomorphy. Using the operator norm on $H^{2}(D_{3/2})$ and Lemma 4.3,
$\|\psi_{n}\,(f\circ\tau_{n})\|_{H^{2}(D_{3/2})}\leq\sup_{w\in D_{3/2}}|\psi_{n}(w)|\ \|f\circ\tau_{n}\|_{H^{2}(D_{3/2})}\leq(n-\tfrac{1}{2})^{-2}\,\|C_{\tau_{n}}\|\,\|f\|_{H^{2}(D_{1})}.$
Summing in $n$ yields
$\|\mathcal{L}f\|_{H^{2}(D_{3/2})}\leq\left(\sum_{n\geq 1}\frac{\|C_{\tau_{n}}\|}{(n-\tfrac{1}{2})^{2}}\right)\|f\|_{H^{2}(D_{1})},$
so $\mathcal{L}$ is bounded and the first inequality follows.
For the explicit constant, combine $\|C_{\tau_{n}}\|\leq\sqrt{2n+1}$ (Lemma 4.5 below) with the comparison $\sqrt{2n+1}\leq\sqrt{6}\,(n-\tfrac{1}{2})^{1/2}$ and the identity $\sum_{n\geq 1}(n-\tfrac{1}{2})^{-3/2}=(2^{3/2}-1)\zeta(3/2)$. ∎
Lemma 4.5.
For every $n\geq 1$, the composition operator
$C_{\tau_{n}}:H^{2}(D_{1})\to H^{2}(D_{3/2}),\qquad C_{\tau_{n}}f=f\circ\tau_{n},$
satisfies
$\|C_{\tau_{n}}\|_{H^{2}(D_{1})\to H^{2}(D_{3/2})}\ \leq\ \sqrt{2n+1}.$
Proof.
Let $T_{r}(\zeta)=1+r\zeta$ map $\mathbb{D}$ biholomorphically onto $D_{r}$, and define the unitary $U_{r}:H^{2}(D_{r})\to H^{2}(\mathbb{D})$ by $(U_{r}f)(\zeta)=f(T_{r}(\zeta))$. Then $U_{3/2}\,C_{\tau_{n}}\,U_{1}^{-1}=C_{\phi_{n}}$ with
$\phi_{n}:=T_{1}^{-1}\circ\tau_{n}\circ T_{3/2}.$
A direct computation gives $\phi_{n}(0)=-\frac{n}{n+1}$. By the standard Hardy-space estimate for composition operators on $H^{2}(\mathbb{D})$ [20],
$\|C_{\phi}\|_{H^{2}(\mathbb{D})}\leq\left(\frac{1+|\phi(0)|}{1-|\phi(0)|}\right)^{1/2},$
hence
$\|C_{\tau_{n}}\|=\|C_{\phi_{n}}\|\leq\left(\frac{1+\frac{n}{n+1}}{1-\frac{n}{n+1}}\right)^{1/2}=\sqrt{2n+1}.$
Let $S:=J_{3/2\to 1}:H^{2}(D_{3/2})\to H^{2}(D_{1})$ denote the restriction map (Definition 4.1). Define
$L_{1}:=S\,\mathcal{L}:H^{2}(D_{1})\to H^{2}(D_{1}),\qquad L_{3/2}:=\mathcal{L}\,S:H^{2}(D_{3/2})\to H^{2}(D_{3/2}).$
Remark 4.6.
Throughout this section we use the subscripted notation $L_{1}$ and $L_{3/2}$ to distinguish the two realizations. From Section 5 onward, we write simply $L$ for the operator $L_{1}=S\,\mathcal{L}$ on $H^{2}(D_{1})$ and $L_{K}$ for its finite-rank truncation.
Proposition 4.7.
The operators $L_{1}$ and $L_{3/2}$ are compact.
Proof.
We will show that $L_{1}$ is the operator-norm limit of finite-rank truncations, and the argument for $L_{3/2}$ is identical.
Let $\Pi_{K}:H^{2}(D_{3/2})\to H^{2}(D_{3/2})$ be the degree-$K$ Taylor truncation at $1$:
$\Pi_{K}\!\left(\sum_{m\geq 0}a_{m}(w-1)^{m}\right)=\sum_{m=0}^{K}a_{m}(w-1)^{m}.$
Then $\Pi_{K}$ has rank $K+1$. Define $(L_{1})_{K}:=\Pi_{K}\,S\,\mathcal{L}$. Since $\Pi_{K}$ is finite rank, so is $(L_{1})_{K}$. In Subsection 4.3 below we prove the tail estimate $\|L_{1}-(L_{1})_{K}\|\leq C_{2}(2/3)^{K+1}\to 0$, which implies that $L_{1}$ is compact. The same construction on the domain side yields compactness of $L_{3/2}$. ∎
4.3 Explicit truncation bounds and construction of $L_{N}$ with $\|L-L_{N}\|\leq\varepsilon_{N}$
Definition 4.8.
For $r>0$ and $K\geq 0$, let $\Pi_{K}:H^{2}(D_{r})\to H^{2}(D_{r})$ be the degree-$K$ truncation at $1$:
$\Pi_{K}f(w):=\sum_{m=0}^{K}c_{m}(w-1)^{m},\qquad f(w)=\sum_{m\geq 0}c_{m}(w-1)^{m}.$
Then $\Pi_{K}$ has rank $K+1$ and operator norm $\|\Pi_{K}\|=1$.
Lemma 4.9.
Let $0<r<R$ and let $J_{R\to r}:H^{2}(D_{R})\to H^{2}(D_{r})$ be restriction. Then for every $K\geq 0$ and every $f\in H^{2}(D_{R})$,
$\|\,J_{R\to r}f-\Pi_{K}(J_{R\to r}f)\,\|_{H^{2}(D_{r})}\ \leq\ \left(\frac{r}{R}\right)^{K+1}\,\|f\|_{H^{2}(D_{R})}.$
In particular, for $(r,R)=(1,3/2)$,
$\|\,J_{3/2\to 1}f-\Pi_{K}(J_{3/2\to 1}f)\,\|_{H^{2}(D_{1})}\ \leq\ \left(\frac{2}{3}\right)^{K+1}\,\|f\|_{H^{2}(D_{3/2})}.$
Proof.
Write $f(w)=\sum_{m\geq 0}c_{m}(w-1)^{m}$ on $D_{R}$. Then
$\|J_{R\to r}f-\Pi_{K}(J_{R\to r}f)\|_{H^{2}(D_{r})}^{2}=\sum_{m>K}|c_{m}|^{2}r^{2m}=\sum_{m>K}|c_{m}|^{2}R^{2m}\left(\frac{r}{R}\right)^{2m}.$
Since $m>K$ implies $(r/R)^{2m}\leq(r/R)^{2(K+1)}$, we obtain
$\|J_{R\to r}f-\Pi_{K}(J_{R\to r}f)\|_{H^{2}(D_{r})}^{2}\leq\left(\frac{r}{R}\right)^{2(K+1)}\sum_{m>K}|c_{m}|^{2}R^{2m}\leq\left(\frac{r}{R}\right)^{2(K+1)}\|f\|_{H^{2}(D_{R})}^{2},$
and taking square roots gives the claim. ∎
Theorem 4.10.
Let $\mathcal{L}:H^{2}(D_{1})\to H^{2}(D_{3/2})$ be the GKW operator from (14), and set
$L_{1}:=S\,\mathcal{L}:H^{2}(D_{1})\to H^{2}(D_{1}),\qquad L_{3/2}:=\mathcal{L}\,S:H^{2}(D_{3/2})\to H^{2}(D_{3/2}).$
For $K,k\geq 0$ define the finite-rank truncations
$(L_{1})_{K}:=\Pi_{K}\,S\,\mathcal{L},\qquad(L_{3/2})_{k}:=\mathcal{L}\,\Pi_{k}\,S.$
Then
$\|L_{1}-(L_{1})_{K}\|_{H^{2}(D_{1})\to H^{2}(D_{1})}\ \leq\ C_{2}\left(\frac{2}{3}\right)^{K+1},\qquad\|L_{3/2}-(L_{3/2})_{k}\|_{H^{2}(D_{3/2})\to H^{2}(D_{3/2})}\ \leq\ C_{2}\left(\frac{2}{3}\right)^{k+1},$
where $C_{2}:=\|\mathcal{L}\|_{H^{2}(D_{1})\to H^{2}(D_{3/2})}$ (cf. Proposition 4.4).
Proof.
(i) Let $f\in H^{2}(D_{1})$ and set $g:=\mathcal{L}f\in H^{2}(D_{3/2})$. Then $L_{1}f=Sg=J_{3/2\to 1}g$. By Lemma 4.9 with $(r,R)=(1,3/2)$,
$\|L_{1}f-(L_{1})_{K}f\|_{H^{2}(D_{1})}=\|\,Sg-\Pi_{K}Sg\,\|_{H^{2}(D_{1})}\leq\left(\frac{2}{3}\right)^{K+1}\|g\|_{H^{2}(D_{3/2})}\leq C_{2}\left(\frac{2}{3}\right)^{K+1}\|f\|_{H^{2}(D_{1})}.$
Taking the supremum over $\|f\|_{H^{2}(D_{1})}=1$ gives the first bound.
(ii) Let $F\in H^{2}(D_{3/2})$. Then
$\|L_{3/2}F-(L_{3/2})_{k}F\|_{H^{2}(D_{3/2})}=\|\mathcal{L}(SF-\Pi_{k}SF)\|_{H^{2}(D_{3/2})}\leq C_{2}\,\|SF-\Pi_{k}SF\|_{H^{2}(D_{1})}.$
Apply Lemma 4.9 with $(r,R)=(1,3/2)$ to $F$ to get $\|SF-\Pi_{k}SF\|_{H^{2}(D_{1})}\leq(2/3)^{k+1}\|F\|_{H^{2}(D_{3/2})}$, yielding the second bound. ∎
Corollary 4.11.
The approximation numbers of $L_{1}$ and $L_{3/2}$ satisfy
$a_{m}(L_{1})\ \leq\ C_{2}\Bigl(\tfrac{2}{3}\Bigr)^{m-1},\qquad a_{m}(L_{3/2})\ \leq\ C_{2}\Bigl(\tfrac{2}{3}\Bigr)^{m-1}\qquad(m\geq 1).$
Hence both operators are trace class, with trace norm ($\|\cdot\|_{S_{1}}:=\sum_{m\geq 1}s_{m}$)
$\|L_{1}\|_{S_{1}}\ \leq\ 3\,C_{2},\qquad\|L_{3/2}\|_{S_{1}}\ \leq\ 3\,C_{2},$
and their eigenvalues $\{\lambda_{m}\}_{m\geq 1}$ (ordered by decreasing modulus, with multiplicity) satisfy
$|\lambda_{m}(L_{1})|,\ |\lambda_{m}(L_{3/2})|\ \leq\ C_{2}\Bigl(\tfrac{2}{3}\Bigr)^{m-1},\qquad\sum_{m\geq 1}|\lambda_{m}|<\infty.$
Proof.
For $m\geq 2$, choose K=m2K=m-2 in Theorem 4.10; then $(L_{1})_{K}$ has rank $\leq m-1$, so
$a_{m}(L_{1})\ \leq\ \|L_{1}-(L_{1})_{m-2}\|\ \leq\ C_{2}\Bigl(\tfrac{2}{3}\Bigr)^{m-1}.$
For $m=1$ we have $a_{1}(L_{1})\leq\|L_{1}\|\leq C_{2}$. On $H^{2}(D_{r})$ the approximation numbers coincide with the singular values, and Weyl’s inequality gives $|\lambda_{m}|\leq s_{m}=a_{m}$ [51, Thm. 1.33]. Finally, $\sum_{m\geq 1}s_{m}(L_{1})\leq C_{2}\sum_{m\geq 1}(2/3)^{m-1}=3C_{2}<\infty$, so $L_{1}$ is trace class [51, Thm. 1.35], and $\sum|\lambda_{m}|\leq\sum s_{m}<\infty$. The argument for $L_{3/2}$ is identical. ∎
5 Certified spectral enclosures for the Gauss–Kuzmin–Wirsing operator
Throughout, the cross-space operator norm
$C_{2}:=\|\mathcal{L}\|_{H^{2}(D_{1})\to H^{2}(D_{3/2})}$
enters only through the explicit truncation budget
$\varepsilon_{K}=C_{2}\Big(\frac{2}{3}\Big)^{K+1},$
so improving the numerical bound for $C_{2}$ immediately sharpens all a posteriori certificates.
5.1 Refined bounds for $C_{2}$ and truncation budgets
Remark 5.1.
Write
$\frac{\sqrt{2n+1}}{(n-\tfrac{1}{2})^{2}}=\sqrt{\frac{2n+1}{\,n-\tfrac{1}{2}\,}}\,(n-\tfrac{1}{2})^{-3/2}.$
For any $N\geq 1$ we have
$\sup_{n\geq N}\frac{2n+1}{n-\tfrac{1}{2}}=2+\frac{2}{N-\tfrac{1}{2}},$
hence, splitting the series at $N$ yields
$C_{2}=\|\mathcal{L}\|_{H^{2}(D_{1})\to H^{2}(D_{3/2})}\leq\sum_{n=1}^{N-1}\frac{\sqrt{2n+1}}{(n-\tfrac{1}{2})^{2}}\;+\;\sqrt{\,2+\frac{2}{N-\tfrac{1}{2}}\,}\,\sum_{n\geq N}(n-\tfrac{1}{2})^{-3/2}.$
The tail can be expressed via the Hurwitz zeta function:
$\sum_{n\geq N}(n-\tfrac{1}{2})^{-3/2}=\zeta\!\left(\tfrac{3}{2},\,N-\tfrac{1}{2}\right).$
Table 1 reports rigorous upper bounds for $C_{2}$ obtained from this splitting strategy, computed with Arb; the maximum enclosure radius is bounded by $5\times 10^{-34}$. (The case $N=1$ coincides with the base estimate from Proposition 4.4.)
NN $C_{2}(N)$
1 11.7000807086
2 10.4849507042
3 10.2709840576
4 10.1907129342
5 10.1506945125
10 10.0893118365
100 10.0590444185
1000 10.0581271764
10000 10.0580982915
Table 1: Refined upper bounds for $C_{2}$ as a function of the splitting NN.
Once a bound for $C_{2}$ is fixed, all truncation budgets used below are explicit and of the form
$\varepsilon_{K}=C_{2}\Big(\frac{2}{3}\Big)^{K+1},$
as in Theorem 4.10. This is the only infinite-dimensional input to the certification step.
Remark 5.2.
Note that as soon as we have a validated bound on $\|L_{K}\|$ we can update $C_{2}$. In particular, if we are able to rigorously prove that $\|L_{K}\|\leq\tilde{C}_{2}$, we have that
$\|L\|\leq\tilde{C}_{2}+\|L\|\Big(\frac{2}{3}\Big)^{K+1}$
and we get the finer bound
$\|L\|\leq\frac{\tilde{C}_{2}}{1-\Big(\frac{2}{3}\Big)^{K+1}}$
5.2 Discretization and matrix assembly
We discretize $L=S\mathcal{L}$ on $H^{2}(D_{1})$ by truncating to the shifted monomial subspace
$\mathcal{V}_{K+1}:=\mathrm{span}\{(w-1)^{j}\}_{j=0}^{K}.$
Let $\Pi_{K}$ denote the corresponding truncation operator (degree $\leq K$). The finite-rank discretization is then represented, in this basis, by a matrix $A_{K}\in\mathbb{C}^{(K+1)\times(K+1)}$.
We assemble $A_{K}$ using exact coefficient formulas expressed in terms of Hurwitz zeta values, evaluated in ball arithmetic (Arb). Concretely, for each column $k$ we expand the image of the basis element $\phi_{k}(w)=(w-1)^{k}$ in a Taylor series at $w=1$,
$(L\phi_{k})(w)=\sum_{\ell\geq 0}(A_{K})_{\ell,k}\,(w-1)^{\ell},\qquad 0\leq k\leq K,$
and retain the coefficients $\ell=0,\dots,K$. The resulting matrix entries are complex balls, hence already include rigorous enclosures for all evaluation/rounding errors at assembly time.
The operator-level truncation error satisfies
$\|L-L_{K}\|\leq\varepsilon_{K},$
with $\varepsilon_{K}$ as above; this is the quantity used in the resolvent perturbation bound and projector perturbation bounds. The numerical assembly and certification results are presented in the following subsections.
5.3 Certified resolvent bounds on contours
To lift the finite-dimensional spectral data of $L_{K}$ to the infinite-dimensional operator $L:H^{2}(D_{1})\to H^{2}(D_{1})$, we verify the small-gain condition: for each eigenvalue $\lambda_{j}$ and an excluding circle $\Gamma_{j}=\{z:|z-\hat{\lambda}_{j}|=r_{j}\}$,
$\alpha_{j}\;=\;\varepsilon_{K}\cdot\sup_{z\in\Gamma_{j}}\|(zI-L_{K})^{-1}\|\;<\;1,$ (15)
where $\varepsilon_{K}=C_{2}(2/3)^{K+1}$ is the truncation error bound (Theorem 4.10). When (15) holds, $\Gamma_{j}\subset\rho(L)$ and $L$ has exactly as many eigenvalues (counted with multiplicity) inside $\Gamma_{j}$ as $L_{K}$.
We employ a two-stage strategy, following the certification workflow of §3.3. For eigenvalues $j=1,\ldots,15$, the resolvent of the full matrix $L_{48}$ is sampled at $256$ points on $\Gamma_{j}$ and certified via Theorem 3.4 (Schur defects $\to$ resolvent bound), yielding $\alpha_{j}\leq 6.2\times 10^{-1}$ ($\varepsilon_{48}\approx 2.37\times 10^{-8}$). For $j=16,\ldots,50$, the eigenvalue gaps shrink below $10^{-7}$ and the resolvent of $L_{48}$ becomes too large; instead, we certify the resolvent at $K=256$ ($\varepsilon_{256}\approx 5.59\times 10^{-45}$) using the block Schur method: the certified Schur form $L_{256}=QTQ^{*}$ is partitioned around $\lambda_{j}$ (Lemma 3.6), and the resolvent supremum on each circle $\Gamma_{j}$ is certified by Proposition 3.8, giving $\alpha_{j}\leq 3.14\times 10^{-4}$ for all $50$ eigenvalues.
The certified finite-dimensional resolvent $\|R_{L_{K}}\|$ is then lifted to the infinite-dimensional one via Neumann perturbation:
$\|R_{L}(z)\|\;\leq\;\frac{\|R_{L_{K}}(z)\|}{1-\alpha_{j}}\;=:\;M_{\infty,j}.$
The full results are given in the supplementary material (Tables S3 and S4); all certified data are also available in machine-readable form in [47]. All $50$ excluding circles are certified, proving that the first $50$ eigenvalues of $L$ are simple.
5.4 Eigenvalue, eigenvector, and projector enclosures
The resolvent certification also yields bounds on the Riesz projector errors. Writing $P_{j}=P_{L}(\Gamma_{j})$ and $P_{K,j}=P_{L_{K}}(\Gamma_{j})$ for the spectral projectors of $L$ and $L_{K}$ associated with the excluding circle $\Gamma_{j}$, the contour integral representation gives
$\vartheta_{j}\;:=\;\|P_{j}-P_{K,j}\|\;\leq\;\frac{|\Gamma_{j}|}{2\pi}\cdot\frac{M_{\infty,j}^{2}\cdot\varepsilon_{K}}{1-\alpha_{j}}.$ (16)
At the finite-dimensional level, the projector error $\|P_{L_{K}}(\Gamma_{j})-\tilde{P}_{T}(\Gamma_{j})\|$ arising from the numerical Schur decomposition and reordering is bounded by the decomposition of Lemma 3.11: the Schur residual contributes (12) and the ordschur reordering contributes (13). At $K=256$, $\vartheta_{j}$ ranges from $1.01\times 10^{-41}$ ($j=1$) to $2.96\times 10^{-3}$ ($j=35$). For $j\geq 36$, useful projector bounds require higher $K$; we use $K=1024$ (see supplementary Table S3).
Since each $\lambda_{j}$ is certified simple and $\vartheta_{j}<1$, Lemma 2.12 gives the eigenvalue error bound
$|\lambda_{j}-\hat{\lambda}_{j}|\;\leq\;\frac{\varepsilon_{K}(1+\vartheta_{j})+2C\,\vartheta_{j}}{1-\vartheta_{j}},$ (17)
where $C=\|L_{K}\|$ is a certified upper bound on the discretization norm and $\varepsilon_{K}=\|L-L_{K}\|$. This avoids any Newton–Kantorovich argument: the projector error alone controls the eigenvalue error for all $50$ eigenvalues simultaneously, with no requirement on the eigenvalue gap relative to $\varepsilon_{K}$. For the dominant eigenvalues ($j\leq 15$) the bound (17) is of order $\varepsilon_{K}\sim 10^{-8}$ (at $K=48$); for $j\leq 35$ at $K=256$ it reaches $\sim 10^{-45}$. The tightest enclosures use $K=1024$ (2048-bit precision, $\varepsilon_{1024}\approx 3.23\times 10^{-180}$): for the leading eigenvalues ($j\leq 15$) the enclosure radii are of order $10^{-176}$$10^{-171}$, degrading to $\sim 10^{-119}$ for $j=50$ due to resolvent growth on the excluding circles. In particular, $|\lambda_{2}-\tilde{\lambda}_{2}|<5\times 10^{-176}$ with $\tilde{\lambda}_{2}$ as in Theorem 1.1, certifying the Wirsing constant to over $175$ decimal digits. The full list of eigenvalue enclosures is given in the supplementary material (Table S3).
The eigenvectors $\hat{v}_{j}$ of $L_{K}$ are the first columns of the ordered Schur unitary $Q_{\mathrm{ord}}$, known to within the combined Schur and ordschur similarity defect $\|E\|\leq 5.72\times 10^{-317}$ at $K=1024$ (supplementary Table S6). At the infinite-dimensional level, the projector error $\vartheta_{j}$ controls the gap between the true eigenvector $v_{j}$ of $L$ and $\hat{v}_{j}$ of $L_{K}$: $\|v_{j}-\hat{v}_{j}\|\leq 2\vartheta_{j}/(1-\vartheta_{j})$.
Since $L$ is non-normal, the spectral expansion requires the dual eigenvectors $\ell_{j}$. We compute the spectral coefficient $\ell_{j}(\mathbf{1})=e_{1}^{*}P_{j}e_{1}$ from the ordered Schur form: with $T_{\mathrm{ord}}=\bigl[\begin{smallmatrix}\lambda_{j}&T_{12}\\ 0&T_{22}\end{smallmatrix}\bigr]$ and $q=Q_{\mathrm{ord}}^{*}e_{1}$,
$\ell_{j}(\mathbf{1})=q_{1}-T_{12}\cdot(T_{22}-\lambda_{j}I)^{-1}q_{\mathrm{rest}},$ (18)
where the triangular system is solved in $P$-bit BigFloat arithmetic. The certified error is dominated by the Schur similarity defect $\|E\|\leq r_{\mathrm{sch}}\sqrt{1+\delta}+C_{A}\,\delta$ (Lemma 3.1); the triangular solve residual is negligible ($\sim 10^{-617}$ at 2048-bit).
At $K=1024$ (2048-bit precision, $\|E\|\leq 5.72\times 10^{-317}$), all $50$ spectral coefficients are sign-certified: $|\ell_{j}(\mathbf{1})|>\delta_{j}$ with radii $\delta_{j}\in[4.5\times 10^{-316},\;1.1\times 10^{-282}]$. The signs alternate as $\ell_{1}>0$, $\ell_{2}>0$, $\ell_{3}<0$, $\ell_{4}>0$, …, with $|\ell_{j}(\mathbf{1})|\sim|\lambda_{j}|$ for large $j$. Cross-validation at $K=256$, $512$, and $1024$ confirms agreement of all certified digits (see supplementary Table S6).
5.5 Certified spectral expansion and computation of $F_{n}(x)$
With all $50$ eigenvalues certified simple and the spectral coefficients $\ell_{j}(\mathbf{1})$ sign-certified, the spectral expansion
$L^{n}\mathbf{1}=\sum_{j=1}^{N}\lambda_{j}^{n}\,\ell_{j}(\mathbf{1})\,v_{j}+R_{N}(n)$ (19)
holds rigorously in $H^{2}(D_{1})$ with the tail bound
$\|R_{N}(n)\|\;\leq\;\rho^{n+1}\cdot M_{\infty}\cdot\|Q_{N}\mathbf{1}\|,$ (20)
where $Q_{N}=I-\sum_{j=1}^{N}P_{j}$ is the tail projector, $\rho$ is the radius of a separating circle in the eigenvalue gap $|\lambda_{N+1}|<\rho<|\lambda_{N}|$, and $M_{\infty}$ is the certified resolvent on this circle. For $N=50$, the parameters are $\rho=1.014\times 10^{-21}$, $M_{\infty}=6.062\times 10^{41}$, and $\|Q_{50}\mathbf{1}\|\leq 6.138\times 10^{-21}$ (computed at $K=1024$), giving $\|R_{50}(n)\|\leq 3.72\times 10^{21}\cdot(1.01\times 10^{-21})^{n+1}$. Already at $n=1$ the remainder is below $10^{-20}$; at $n=5$ it drops below $10^{-104}$ (see supplementary Table S8).
Define the integrated spectral expansion
$F_{n}^{(j_{0})}(x)\;=\;\int_{0}^{x}\sum_{j=j_{0}}^{N}\lambda_{j}^{n}\,\ell_{j}(\mathbf{1})\,v_{j}(t)\,dt,$
which is computed exactly in the shifted monomial basis via $\int_{0}^{x}(t-1)^{k}\,dt=\bigl((x{-}1)^{k+1}-(-1)^{k+1}\bigr)/(k{+}1)$. For $j_{0}=2$, this measures the integrated deviation of $L^{n}\mathbf{1}$ from its equilibrium $\ell_{1}(\mathbf{1})\,v_{1}$. The $L^{\infty}$ error is controlled by Cauchy–Schwarz and the Gram matrix of the basis on $[0,1]$ (spectral norm $\leq\pi$):
$\sup_{x\in[0,1]}\Bigl|F_{n}^{(j_{0})}(x)-\int_{0}^{x}\!\Bigl(L^{n}\mathbf{1}-\textstyle\sum_{j<j_{0}}\lambda_{j}^{n}P_{j}\mathbf{1}\Bigr)dt\Bigr|\;\leq\;\sqrt{\pi}\;\rho^{n+1}\cdot C\;\leq\;6.6\times 10^{21}\cdot(1.01\times 10^{-21})^{n+1}.$
Figure 2 shows $F_{n}^{(j_{0})}$ for $j_{0}=2,\ldots,6$ and $n=1,2,5$. As dominant eigenfunctions are progressively removed, the Markov partition structure of the Gauss map becomes increasingly visible, and the amplitude decreases by roughly a factor $|\lambda_{j_{0}}/\lambda_{j_{0}-1}|\approx 1/3$ per removed mode.
Figure 2: Integrated spectral expansion $F_{n}^{(j_{0})}(x)=\int_{0}^{x}\sum_{j\geq j_{0}}\lambda_{j}^{n}\,\ell_{j}(\mathbf{1})\,v_{j}(t)\,dt$ with the first $j_{0}-1$ eigenfunctions cumulatively removed. Each panel shows $n=1$ (blue), $n=2$ (red), $n=5$ (green). Dotted vertical lines mark the Markov partition boundaries $x=1/k$. For $j_{0}=6$, $n=5$, the signal amplitude is  ${\sim}\,10^{-16}$ while the $L^{\infty}$ error bound is  ${\leq}\,10^{-105}$, confirming that the plotted curves are rigorous to plotting accuracy.
Returning to the Gauss–Kuzmin problem, the distribution function $G_{n}(x)=\int_{0}^{x}L^{n}\mathbf{1}(t)\,dt$ converges to $\log_{2}(1+x)$. Since $\lambda_{1}=1$ and $\ell_{1}(\mathbf{1})\,v_{1}=\frac{1}{\ln 2}\cdot\frac{1}{1+x}$ (the invariant density), the spectral expansion gives
$|G_{n}(x)-\log_{2}(1{+}x)|\;\leq\;|F_{n}^{(2)}(x)|+\sqrt{\pi}\;\|R_{N}(n)\|.$
The dominant error term is $\lambda_{2}^{n}\,\ell_{2}(\mathbf{1})\,\int_{0}^{x}v_{2}$, which decays at rate $|\lambda_{2}|^{n}\approx 0.3037^{n}$, the Wirsing constant. The tail bound certifies that the $50$-term expansion captures this rate with rigorous error below $10^{-20}$ already at $n=1$. Figure 3 illustrates the convergence: the spectral approximant $S_{50}(n,x)/\lambda_{1}^{n}$ rapidly approaches the invariant density $\ell_{1}(\mathbf{1})\,v_{1}(x)$, with the spectral gap $|\lambda_{2}/\lambda_{1}|\approx 0.304$ governing the rate.
Figure 3: Convergence of the spectral approximant $S_{50}(n,x)=\sum_{j=1}^{50}\lambda_{j}^{n}\,\ell_{j}(\mathbf{1})\,v_{j}(x)$ (normalized by $\lambda_{1}^{n}$) to the invariant density $\ell_{1}(\mathbf{1})\,v_{1}(x)=\frac{1}{\ln 2(1+x)}$ (dashed black). The spectral gap $|\lambda_{2}/\lambda_{1}|\approx 0.304$ governs the convergence rate: by $n=5$ the approximation is indistinguishable from the limit.
6 Discussion and portability
Moving the bottleneck from analysis to certification.
The central contribution is a shift in where the hard work lives. Once an operator $L$ admits compactness/nuclearity on a suitable space or a DFLY inequality (Appendix A), the remaining spectral analysis reduces to checkable numerical statements about a finite-rank model $L_{N}$: a certified approximation bound $\|L-L_{N}\|\leq\varepsilon_{N}$, certified resolvent bounds along chosen contours, and certified residual information from linear algebra. Theorems are phrased in a fully a posteriori form, and a large portion of the proof burden is converted into validated numerical linear algebra, building on [12]. A practical consequence is high-throughput certification: many eigenvalues, multiplicities, and Riesz projectors can be validated simultaneously, precisely in the regime where non-normality and pseudospectral effects make resolvent-based certification essential [31].
The spectral-gap bottleneck and a posteriori reversal.
Since a spectral gap is often the hardest ingredient in rigorous thermodynamic computations [48, 49, 55], the a posteriori reversal described in Section 1 is of particular interest. Liverani [43] proposed certifying spectral data of the true operator from a finite-rank discretization via Keller–Liverani perturbation theory [36], but the approach requires the strong resolvent norm of the discretization, which is difficult to compute when the strong norm is, e.g., the BV norm. In the DFLY setting of Appendix A, we circumvent this by working with the weak resolvent of the discretization, which is computable via certified singular value enclosures of a finite matrix, and lifting to strong resolvent bounds via the Lasota–Yorke inequality (Proposition A.7; see Remark A.8 for a detailed comparison). This makes the a posteriori approach both complementary to a priori methods on holomorphic function spaces [6, 7] or structured approximation schemes [8, 9, 52], and fully implementable.
Full spectral convergence in the DFLY setting.
A second main contribution, developed in Appendix A, is the convergence guarantee of Theorem A.18: under a DFLY inequality and a mild rate condition on the discretization error, every isolated eigenvalue, eigenvector, and Riesz projector outside the essential spectral radius can be approximated to arbitrary precision by the finite-rank data of $L_{k}$. This upgrades the qualitative spectral stability guaranteed by the Keller–Liverani theorem [36] (which itself extends Li’s Ulam-conjecture resolution [42] beyond the invariant density) to a fully certified and quantitative statement. A key ingredient is the coarse-to-fine propagation of Proposition A.14, which controls the growth of the weak resolvent $\mathcal{R}_{w}(z,L_{k})$ as $k\to\infty$, preventing the spectral pollution that would otherwise invalidate finer certifications.
Portability.
[…]
Limitations.
The method’s quantitative strength is governed by the product $\varepsilon_{N}\cdot\sup_{z\in\Gamma}\|(zI-L_{N})^{-1}\|$: if this exceeds $1$, the Neumann-series bound fails; if it is close to $1$, constants become weak, reflecting genuine pseudospectral sensitivity. The pipeline certifies isolated spectral data (clusters separated by contours), not continuous spectrum.
Reproducibility.
All certifications were carried out on a single desktop workstation with no supercomputing resources (see supplementary material for specifications); increasing the number of eigenvalues or the precision scales directly with matrix size and arithmetic precision. Detailed computational methodology, all certified spectral tables, eigenvector coefficients, and additional figures are collected in the supplementary material (arXiv ancillary file). The full source code for the certification pipeline is available at https://github.com/orkolorko/GKWExperiments.jl; certified spectral data in machine-readable form are deposited at https://doi.org/10.7910/DVN/HKM3Y2.
Appendix A Strong–weak (DFLY) resolvent lifting
This appendix extends the abstract compact template of Section 2 to the common strong-weak transfer-operator setting: a Banach scale $(\mathcal{B}_{s},\|\cdot\|_{s})\hookrightarrow(\mathcal{B}_{w},\|\cdot\|_{w})$ with compact embedding, and a Doeblin–Fortet–Lasota–Yorke (DFLY) inequality. The same resolvent-based certification logic applies: validated weak resolvent bounds for finite-rank discretizations yield certified strong resolvent bounds for the true operator, certified spectral exclusion curves, and a coarse-fine workflow in which a coarse certificate propagates to arbitrarily fine levels.
We use the resolvent notation of Section 2.1, writing $\mathcal{R}_{s}(z,A):=\|R_{A}(z)\|_{s\to s}$ and $\mathcal{R}_{w}(z,A):=\|R_{A}(z)\|_{w\to w}$ to distinguish operator norms.
A.1 Setting: Banach scale, DFLY bounds, and Galerkin hierarchy
Assumption A.1.
Let $(\mathcal{B}_{s},\|\cdot\|_{s})\hookrightarrow(\mathcal{B}_{w},\|\cdot\|_{w})$ be a continuous embedding with embedding constant $E_{s\to w}:=\|i\|_{s\to w}$. Assume the embedding is compact.
Assumption A.2.
Let $L\in\mathcal{L}(\mathcal{B}_{s})$ admit a bounded extension to $\mathcal{B}_{w}$ satisfying the one-step DFLY bounds: for constants $a,b,M\in(0,\infty)$ and all $u\in\mathcal{B}_{s}$,
$\|Lu\|_{s}\leq a\|u\|_{s}+b\|u\|_{w},\qquad\|Lu\|_{w}\leq M\|u\|_{w}.$ (21)
Assume we are given finite-rank operators $L_{k}\in\mathcal{L}(\mathcal{B}_{s})$ extending boundedly to $\mathcal{B}_{w}$ with the same one-step bounds:
$\|L_{k}u\|_{s}\leq a\|u\|_{s}+b\|u\|_{w},\qquad\|L_{k}u\|_{w}\leq M\|u\|_{w}.$ (22)
Define the mixed discrepancy $\delta_{k}:=\|L-L_{k}\|_{s\to w}$.
Assumption A.3.
Assume $L_{k}$ arise from a nested Galerkin hierarchy: finite-rank projections $\pi_{k}:\mathcal{B}_{w}\to\mathcal{B}_{w}$ with $\pi_{k}^{2}=\pi_{k}$, $\|\pi_{k}\|_{w\to w}\leq 1$, $\|\pi_{k}\|_{s\to s}\leq 1$, $\pi_{k}\pi_{k^{\prime}}=\pi_{k}$ for $k^{\prime}\geq k$, $\mathrm{Ran}(\pi_{k})\subset\mathcal{B}_{s}$, and $L_{k}:=\pi_{k}\,L\,\pi_{k}$. Assume there exists a computable sequence $\varepsilon_{k}\downarrow 0$ such that
$\|(\mathrm{Id}-\pi_{k})L\|_{s\to w}+\|L(\mathrm{Id}-\pi_{k})\|_{s\to w}\leq\varepsilon_{k}.$ (23)
In particular, $\|L-L_{k}\|_{s\to w}\leq\varepsilon_{k}$ and $\|L_{k^{\prime}}-L_{k}\|_{s\to w}\leq\varepsilon_{k}$ for $k^{\prime}\geq k$.
Lemma A.4.
Under Assumption A.2, for every $n\geq 1$ and $u\in\mathcal{B}_{s}$,
$\|L^{n}u\|_{s}\leq a^{n}\|u\|_{s}+b_{n}\|u\|_{w},\qquad\|L^{n}u\|_{w}\leq M^{n}\|u\|_{w},$ (24)
where $b_{n}:=b\sum_{j=0}^{n-1}a^{\,n-1-j}M^{\,j}$. The same bounds hold with $L$ replaced by $L_{k}$.
Proof.
Iterate the weak bound and insert $\|L^{j}u\|_{w}\leq M^{j}\|u\|_{w}$ into the strong inequality. ∎
Lemma A.5.
Under Assumption A.2, if (zIL)u=f(zI-L)u=f with $u,f\in\mathcal{B}_{s}$, then
$(|z|-a)\,\|u\|_{s}\ \leq\ b\,\|u\|_{w}+\|f\|_{s}.$ (25)
The same holds with $L$ replaced by $L_{k}$.
Proof.
From $zu=Lu+f$ and (21): $|z|\|u\|_{s}\leq a\|u\|_{s}+b\|u\|_{w}+\|f\|_{s}$. ∎
Remark A.6.
Under Assumptions A.1A.2, the essential spectral radius of $L:\mathcal{B}_{s}\to\mathcal{B}_{s}$ is $\leq a$ (Hennion-type quasi-compactness). We use only: for $|z|>a$, injectivity of zILzI-L implies invertibility on $\mathcal{B}_{s}$.
A.2 Weak-to-strong resolvent lifting and spectral exclusion
Proposition A.7.
Assume Assumption A.1A.2. Fix $k\geq 1$ and $z\in\mathbb{C}$ with $z\notin\sigma(L_{k})$, and set $M_{k}(z):=\mathcal{R}_{w}(z,L_{k})$. If
$|z|\ >\ a+b\,M_{k}(z)\,\delta_{k},$ (26)
then $z\notin\sigma(L)$ and
$\mathcal{R}_{s}(z,L)\ \leq\ \frac{b\,M_{k}(z)\,E_{s\to w}+1}{|z|-a-b\,M_{k}(z)\,\delta_{k}}.$ (27)
Proof.
Injectivity. Let (zIL)u=0(zI-L)u=0. Then $(zI-L_{k})u=(L-L_{k})u$, hence $u=R_{L_{k}}(z)(L-L_{k})u$. Taking weak norms gives $\|u\|_{w}\leq M_{k}(z)\delta_{k}\|u\|_{s}$. Lemma A.5 with $f=0$ yields $(|z|-a)\|u\|_{s}\leq b\,M_{k}(z)\delta_{k}\|u\|_{s}$, forcing $u=0$ under (26). Since $|z|>a$, Remark A.6 gives bijectivity.
Quantitative bound. For $f\in\mathcal{B}_{s}$ with $\|f\|_{s}=1$, set $u:=R_{L}(z)f$. From $(zI-L_{k})u=f+(L-L_{k})u$, taking weak norms yields $\|u\|_{w}\leq M_{k}(z)E_{s\to w}+M_{k}(z)\delta_{k}\|u\|_{s}$. Lemma A.5 applied to (zIL)u=f(zI-L)u=f gives $(|z|-a)\|u\|_{s}\leq b\|u\|_{w}+1$; inserting the weak bound and rearranging yields (27). ∎
Remark A.8.
The use of the weak resolvent $\mathcal{R}_{w}(z,L_{k})$ of the discretization in Proposition A.7 is essential for two reasons.
First, it is computable: since $L_{k}$ is finite rank, $\mathcal{R}_{w}(z,L_{k})$ reduces to a certified singular value problem for a finite matrix. An approach based on the strong resolvent $\mathcal{R}_{s}(z,L_{k})$ (as proposed in the feasibility study [43]) requires the induced $\|\cdot\|_{s}$-norm of a matrix, which is nontrivial when $\|\cdot\|_{s}$ is, e.g., the BV norm.
Second, $\mathcal{R}_{w}(z,L)$ for the original operator is not available: in typical DFLY applications $L$ is a weak contraction ($\|L\|_{w\to w}\leq 1$), so its weak essential spectral radius is $1$ and $\mathcal{R}_{w}(z,L)=+\infty$ for every $|z|\leq 1$, precisely where the discrete eigenvalues of interest lie. A strong-weak Neumann-series argument using $\mathcal{R}_{w}(z,L)$ therefore fails at exactly the points it would be needed. The DFLY inequality resolves this: it converts the computable quantity $\mathcal{R}_{w}(z,L_{k})$ (finite because $L_{k}$ has no essential spectrum) into a certified bound on $\mathcal{R}_{s}(z,L)$ (finite because the eigenvalues are discrete in $\mathcal{B}_{s}$), bypassing $\mathcal{R}_{w}(z,L)$ entirely.
Proposition A.9.
Assume Assumption A.1A.2. Let $\Gamma$ be a piecewise $C^{1}$ closed curve with $\Gamma\subset\rho(L_{k})$ and
$\inf_{z\in\Gamma}\bigl(|z|-a\bigr)\;>\;b\,\delta_{k}\,\sup_{z\in\Gamma}\mathcal{R}_{w}(z,L_{k}).$ (28)
Then $\Gamma\subset\rho(L)$ and $\sup_{z\in\Gamma}\mathcal{R}_{s}(z,L)\leq\sup_{z\in\Gamma}\frac{b\,\mathcal{R}_{w}(z,L_{k})\,E_{s\to w}+1}{|z|-a-b\,\mathcal{R}_{w}(z,L_{k})\,\delta_{k}}$. In particular, if $\Gamma=\partial U$ then $P_{L}(U)$ is well-defined on $\mathcal{B}_{s}$.
Proof.
Apply Proposition A.7 pointwise on $\Gamma$. ∎
Lemma A.10.
If $L_{k}$ satisfies (22) and $z\in\rho(L_{k})$ with $|z|>a$, then
$\mathcal{R}_{s}(z,L_{k})\ \leq\ \frac{b\,\mathcal{R}_{w}(z,L_{k})\,E_{s\to w}+1}{|z|-a}.$ (29)
Proof.
Apply Lemma A.5 to $u:=R_{L_{k}}(z)f$ and bound $\|u\|_{w}$ via $\mathcal{R}_{w}(z,L_{k})$. ∎
Lemma A.11.
If $z\in\rho(L)\cap\rho(L_{k})$, then
$\|R_{L_{k}}(z)-R_{L}(z)\|_{s\to w}\ \leq\ \mathcal{R}_{w}(z,L_{k})\,\delta_{k}\,\mathcal{R}_{s}(z,L).$ (30)
Proof.
From $R_{L_{k}}(z)-R_{L}(z)=R_{L_{k}}(z)(L-L_{k})R_{L}(z)$, bound $\|(L-L_{k})v\|_{w}\leq\delta_{k}\|v\|_{s}$. ∎
A.3 Coarse–fine propagation and weak resolvent growth
A key issue in the DFLY framework is to prevent the weak resolvent $\mathcal{R}_{w}(z,L_{k})$ from exploding as $k$ increases. The next results extend the resolvent-blowup control of [12] to the present abstract setting, providing true-to-fine control: once $z\in\rho(L)$ and $\mathcal{R}_{s}(z,L)$ are certified at a coarse level, $\mathcal{R}_{w}(z,L_{k})$ is explicitly controlled for all finer $k$.
Assumption A.12.
For each $k$, set $E_{k}:=\mathrm{Ran}(\pi_{k})$ and assume the norm equivalence constant $E_{k,w\to s}:=\sup_{u\in E_{k}\setminus\{0\}}\|u\|_{s}/\|u\|_{w}$ is explicitly bounded.
Lemma A.13.
Under Assumptions A.2A.12, for every $k$, $N\geq 1$, and $f\in E_{k}$,
$\|L_{k}^{N}f\|_{s}\ \leq\ \bigl(a^{N}E_{k,w\to s}+b_{N}\bigr)\|f\|_{w}.$ (31)
Proof.
Apply Lemma A.4 to $f\in E_{k}\subset\mathcal{B}_{s}$ and use $\|f\|_{s}\leq E_{k,w\to s}\|f\|_{w}$. ∎
Proposition A.14.
Assume Assumption A.1A.3 and Assumption A.12. Fix $k\geq 1$ and $z\in\mathbb{C}$ with $|z|>a$ and $z\in\rho(L)$. Write $K(z):=\mathcal{R}_{s}(z,L)$ and $\Delta_{k}:=\|L-L_{k}\|_{s\to w}$.
Let $N\geq 1$ and define
$S_{N}^{(k)}(z):=\frac{1}{|z|}\sum_{\ell=0}^{N-1}\frac{\|L_{k}^{\ell}\|_{w\to w}}{|z|^{\ell}}\ \leq\ \frac{1}{|z|}\sum_{\ell=0}^{N-1}\Bigl(\frac{M}{|z|}\Bigr)^{\ell},$ (32)
and
$\widetilde{\beta}_{N}(z):=\frac{1}{|z|^{N}}\,\Delta_{k}\,K(z)\,\bigl(a^{N}E_{k,w\to s}+b_{N}\bigr).$ (33)
If $\widetilde{\beta}_{N}(z)<1$, then $z\in\rho(L_{k})$ and
$\mathcal{R}_{w}(z,L_{k})\ \leq\ \frac{S_{N}^{(k)}(z)+\;|z|^{-N}\,E_{s\to w}K(z)\,\bigl(a^{N}E_{k,w\to s}+b_{N}\bigr)}{1-\widetilde{\beta}_{N}(z).}$ (34)
Proof.
Fix $f\in\mathcal{B}_{w}$ and consider $u\in\mathcal{B}_{s}$ solving $(zI-L_{k})u=f$. (Existence will follow from the a priori bound below.) Rewrite the equation as
$(zI-L)u=f+(L-L_{k})u.$
Applying $R_{L}(z)$ yields
$u=R_{L}(z)f+R_{L}(z)(L-L_{k})u.$
Taking weak norms and using $\|(L-L_{k})u\|_{w}\leq\Delta_{k}\|u\|_{s}$ gives
$\|u\|_{w}\leq E_{s\to w}K(z)\|f\|_{s}+E_{s\to w}K(z)\Delta_{k}\|u\|_{s}.$ (35)
Next, apply the truncated Laurent identity: from $(zI-L_{k})u=f$,
$u=\frac{1}{z}\sum_{\ell=0}^{N-1}\frac{L_{k}^{\ell}f}{z^{\ell}}+\frac{1}{z^{N}}\,L_{k}^{N}u.$
Taking weak norms yields
$\|u\|_{w}\leq S_{N}^{(k)}(z)\,\|f\|_{w}+|z|^{-N}\,\|L_{k}^{N}u\|_{w}.$
Using $\|L_{k}^{N}u\|_{w}\leq E_{s\to w}\|L_{k}^{N}u\|_{s}$ and Lemma A.13 (with $L_{k}^{N}$ landing in the finite-dimensional range), we get
$\|L_{k}^{N}u\|_{w}\leq E_{s\to w}\bigl(a^{N}\|u\|_{s}+b_{N}\|u\|_{w}\bigr).$
Insert this into the previous inequality and rearrange to isolate $\|u\|_{w}$:
$\Bigl(1-|z|^{-N}E_{s\to w}b_{N}\Bigr)\|u\|_{w}\leq S_{N}^{(k)}(z)\|f\|_{w}+|z|^{-N}E_{s\to w}a^{N}\|u\|_{s}.$
Now substitute the bound (35) into the right-hand side through $\|u\|_{s}$ and collect terms so that the coefficient of $\|u\|_{w}$ becomes $1-\widetilde{\beta}_{N}(z)$; one obtains
$\|u\|_{w}\leq\frac{S_{N}^{(k)}(z)\|f\|_{w}+|z|^{-N}E_{s\to w}K(z)\bigl(a^{N}E_{k,w\to s}+b_{N}\bigr)\|f\|_{w}}{1-\widetilde{\beta}_{N}(z)},$
where we also used $\|f\|_{s}\leq E_{k,w\to s}\|f\|_{w}$ when $f\in E_{k}$. This proves that $(zI-L_{k})$ is injective with bounded inverse on the finite-dimensional range, hence $z\in\rho(L_{k})$ and (34) holds. ∎
Remark A.15.
In practice one certifies $z\in\rho(L)$ and $K(z)$ on a contour $\Gamma$ using Proposition A.7 at a coarse level $k_{0}$. Proposition A.14 then bounds $\mathcal{R}_{w}(z,L_{k})$ for all finer $k\geq k_{0}$, without recomputing resolvents.
Corollary A.16.
Assume Assumption A.1A.3 and Assumption A.12, with $0<a<M$. Fix $\mu\in(a,M)$ and set
$\mathfrak{q}:=\frac{\bigl|\log(\mu/M)\bigr|}{\bigl|\log(a/M)\bigr|}\in(0,1),\qquad C_{*}:=1+\frac{b}{M-a}.$
Let $k\geq 1$ and assume $z\in\rho(L)$ satisfies $|z|\geq\mu$. Write $K(z):=\mathcal{R}_{s}(z,L)$ and $\delta_{k}:=\|L-L_{k}\|_{s\to w}$. Choose $N_{k}:=\lceil\log E_{k,w\to s}/|\log(a/M)|\rceil$. If
$1-\frac{M}{\mu}\,\delta_{k}\,K(z)\,C_{*}\,E_{k,w\to s}^{\mathfrak{q}}>0,$ (36)
then $z\in\rho(L_{k})$ and
$\mathcal{R}_{w}(z,L_{k})\ \leq\ \frac{\dfrac{M}{\mu}\,E_{k,w\to s}^{\mathfrak{q}}}{1-\dfrac{M}{\mu}\,\delta_{k}\,K(z)\,C_{*}\,E_{k,w\to s}^{\mathfrak{q}}}\left(\frac{1}{M-\mu}+E_{s\to w}\,K(z)\,C_{*}\right).$ (37)
Proof.
Apply Proposition A.14 with $N=N_{k}$. The choice ensures $a^{N_{k}}E_{k,w\to s}\leq M^{N_{k}}$ and $|z|^{-N_{k}}\leq(M/\mu)E_{k,w\to s}^{\mathfrak{q}}$ by logarithmic interpolation. Substituting $b_{N_{k}}\leq\frac{b}{M-a}M^{N_{k}}$ into (34) yields (37). ∎
A.4 Spectral convergence in the DFLY setting
Corollary A.17.
In the setting of Corollary A.16, suppose that at some coarse level $k_{0}$ the perturbation condition (26) certifies $z\in\rho(L)$ and $K(z)<\infty$. If
$\delta_{k}\,E_{k,w\to s}^{\mathfrak{q}}\ \longrightarrow\ 0\qquad\text{as }k\to\infty,$ (38)
then $\|R_{L_{k}}(z)-R_{L}(z)\|_{s\to w}\to 0$: Corollary A.16 gives $\mathcal{R}_{w}(z,L_{k})=O(E_{k,w\to s}^{\mathfrak{q}})$, and Lemma A.11 yields the claim.
The following theorem is the DFLY analogue of Theorem 2.15: it shows that under the feasibility condition (38), the full discrete spectral picture of $L$ outside the essential spectral radius can be recovered from the Galerkin hierarchy $L_{k}$.
Theorem A.18.
Under Assumption A.1A.3 and Assumption A.12, suppose $0<a<M$. Fix $\mu\in(a,M)$ and let $\mathfrak{q}:=|\log(\mu/M)|/|\log(a/M)|\in(0,1)$. Suppose
$\delta_{k}\,E_{k,w\to s}^{\mathfrak{q}}\;\longrightarrow\;0\qquad\text{as }k\to\infty.$ (39)
Let $U\subset\mathbb{C}$ be a bounded domain with piecewise $C^{1}$ boundary $\Gamma:=\partial U$ such that $\inf_{z\in\Gamma}|z|\geq\mu$, $\Gamma\subset\rho(L)$, and $\sigma(L)\cap\Gamma=\emptyset$. Then there exists $k_{0}$ such that for all $k\geq k_{0}$:
  1. 1.
    $\Gamma\subset\rho(L_{k})$ and the perturbation condition $\sup_{z\in\Gamma}b\,\delta_{k}\,\mathcal{R}_{w}(z,L_{k})<\inf_{z\in\Gamma}(|z|-a)$ holds;
  2. 2.
    $\dim\operatorname{Ran}P_{L_{k}}(U)=\dim\operatorname{Ran}P_{L}(U)$;
  3. 3.
    $\|P_{L_{k}}(U)-P_{L}(U)\|_{s\to w}\to 0$ as $k\to\infty$.
In particular, every isolated eigenvalue, eigenvector, and Riesz projector of $L$ outside the essential spectral radius can be approximated to arbitrary precision by the finite-rank data of $L_{k}$, provided the contour can be chosen with $\inf|z|\geq\mu$ for some $\mu$ satisfying (39).
Proof.
Since $\inf_{z\in\Gamma}|z|\geq\mu>a$ and $\Gamma\subset\rho(L)$, Corollary A.16 applies to every $z\in\Gamma$. By Corollary A.17, the rate condition (39) ensures that $\Gamma\subset\rho(L_{k})$ for $k\geq k_{0}$ and $\|R_{L_{k}}(z)-R_{L}(z)\|_{s\to w}\to 0$ uniformly on $\Gamma$. In particular, the perturbation condition of Proposition A.9 is eventually satisfied on all of $\Gamma$, giving (i). Claims (ii) and (iii) follow from the contour integral representation $P_{L_{k}}(U)-P_{L}(U)=\frac{1}{2\pi i}\oint_{\Gamma}(R_{L_{k}}(z)-R_{L}(z))\,dz$ and the uniform convergence of the integrand in $\|\cdot\|_{s\to w}$ established in Corollary A.17. ∎
Acknowledgements
The author would like to acknowledge helpful discussions with A. Blumenthal, O. Bandtlow, and S. Galatolo. The idea of treating the Gauss map arose during a discussion with I. Melbourne. The author is also grateful to the participants of related seminars and workshops for valuable feedback on preliminary versions of this work.
Claude (Anthropic) and ChatGPT (OpenAI) were used as assistants for prose editing and code development.
The author thanks the Uppsala University Library for access to their edition of Gauss’s Werke, the Maurizio Monge Foundation for supporting access to Claude Pro, and Prof. Y. Sato for access to computational resources.
I.N. was partially supported by the Postgraduate Program in Mathematics at UFRJ, CAPES–Finance Code 001, CNPq Projeto Universal No. 404943/2023-3, CAPES–PRINT No. 88881.311616/2018-00, and CAPES–STINT No. 88887.155746/2017-00.
Any remaining errors are solely the author’s responsibility.